### v2

# Dynamic Financial Analysis (DFA) of General Insurers under Climate Change

Abstract
Climate change is expected to significantly affect the physical, financial, and economic environments over the long term, posing risks to the financial health of general insurers. While general insurers typically use Dynamic Financial Analysis (DFA) for a comprehensive view of financial impacts, traditional DFA as presented in the literature does not consider the impact of climate change. To address this gap, we introduce a climate-dependent DFA approach that integrates climate risk into DFA, providing a holistic assessment of the long-term impact of climate change on the general insurance industry.
The proposed framework has three key features. First, it captures the long-term impact of climate change on the assets and liabilities of general insurers by considering both physical and economic dimensions across different climate scenarios within an interconnected structure. Second, it addresses the uncertainty of climate change impacts using stochastic simulations within climate scenario analysis that are useful for actuarial applications. Finally, the framework is tailored to the general insurance sector by addressing its unique characteristics.
To demonstrate the practical application of our model, we conduct an extensive empirical study using Australian data to assess the long-term financial impact of climate change on the general insurance market under various climate scenarios. The results show that the interaction between economic growth and physical risk plays a key role in shaping general insurers’ risk-return profiles. Limitations of our framework are thoroughly discussed.
keywords:
Climate change, Dynamic Financial Analysis, General insurance JEL Codes: Q54 , C53 , G22 MSC classes: 91G70 , 91G60 , 62P05 , 91B30
1 Introduction
1.1 Background
Climate change poses a multifaceted and profound impact on the financial performance and position of general insurers. On the liabilities side, shifting weather patterns are likely to alter the frequency and severity of future claims (see, e.g., Haug et al., 2011; Lyubchich et al., 2019), which could lead to increased claims costs and adversely affect underwriting profitability. On the assets side, climate change-induced fluctuations in temperature, shifting patterns of natural disasters, and evolving socio-economic conditions may significantly influence key macroeconomic variables. These include inflation rates (see, e.g., Parker, 2018; Economides and Xepapadeas, 2018; Kotz et al., 2024), interest rates (see, e.g., Bylund and Jonsson, 2020; Mongelli et al., 2022), and equity returns (see, e.g., Karydas and Xepapadeas, 2022; Venturini, 2022; Barnett, 2023), thereby impacting the investment performance of general insurers. Ultimately, the combined effects of climate change on both assets and liabilities will cascade into insurers’ capital positions, potentially affecting the financial health and stability of the broader general insurance market.
To navigate these challenges, a holistic understanding of climate change’s impact on general insurers is essential. This need is increasingly recognised by regulators, as evidenced by the growing number of climate-related disclosure requirements, such as IFRS S2 (IFRS, 2023) and its Australian equivalent, AASB S2 (AASB, 2024). In the existing literature, several studies have investigated the impacts of climate change on individual financial items. For example, on the liabilities side, various studies have examined climate change effects on hazards such as floods (see, e.g., Seneviratne et al., 2021; Bourget et al., 2024), bushfires (see, e.g., Quilcaille et al., 2022), and tropical cyclones and storms (see, e.g., Jagger et al., 2008, 2011; Meiler et al., 2022), generally indicating that rising temperatures exacerbate the frequency and severity of most hazard events (Chapter 11 of IPCC, 2021a, pp. 1546–1600). Additionally, the impact of climate change on non-catastrophe, weather-related claims frequency and severity have also been investigated in literature (see, e.g., Haug et al., 2011; Lyubchich et al., 2017). Outside the general insurance domain, the impacts of climate change on mortality rates have also been examined in the literature (Min et al., 2025; Guibert et al., 2025). On the assets side, climate change impacts on interest rates, inflation rates, and equity returns have been explored both theoretically using General Equilibrium Models (Economides and Xepapadeas, 2018; Karydas and Xepapadeas, 2022; Barnett, 2023) and empirically using multiple regression (see, e.g., Parker, 2018) and Factor Models (see, e.g., Venturini, 2022). While efforts have been made to incorporate climate change impacts into the modelling of insurance liabilities and assets in various special (and limited) cases, as yet a holistic approach to assess the interconnected financial impacts of climate change on general insurers across both assets and liabilities remains lacking.
A widely used tool in the general insurance industry for informing strategic decisions is Dynamic Financial Analysis (DFA). DFA offers a comprehensive perspective on the financial impacts that affect general insurers (Coutts and Devitt, 1989; Paulson and Dixit, 1989; Kaufmann et al., 2001; Eling and Toplek, 2009). It encompasses a suite of methods designed to project and analyse the future financial position of insurers under various scenarios. Typical applications of DFA include economic capital modelling, solvency monitoring, and strategy testing. Additionally, it functions as a model office, enabling management to evaluate current strategies in a theoretical environment under different scenarios, thereby helping to avoid costly real-world errors (Kaufmann et al., 2001; Eling and Parnitzke, 2007). It can also be employed to analyse market behaviors within the general insurance sector (see, e.g., Taylor, 2008). However, existing DFA frameworks in literature (see, e.g., Kaufmann et al., 2001; D’Arcy and Gorvett, 2004; Consigli et al., 2018) do not explicitly incorporate climatic factors, which limits their effectiveness in addressing the financial implications of climate change. To address this gap, we propose a comprehensive yet tractable “climate-dependent DFA” framework for examining the multifaceted impacts of climate change on the general insurance market, as detailed in the following section.
1.2 Statement of contributions
In this paper, we extend the traditional DFA framework by integrating climate change considerations to provide a holistic assessment of climate impacts on general insurance markets. This framework can serve as an initial step in guiding decision-making for both insurers and regulators facing climate-related challenges. Specifically, we propose the “climate-dependent DFA” framework, designed to address key attributes of climate change impacts within the general insurance context:
  • 1.
    Holistic nature of climate impacts: Climate change can affect both insurance claims and investment returns, impacting insurers’ assets and liabilities. However, most studies assess these effects in isolation, which may misconstrue the overall financial impact due to asset–liability interdependencies. Furthermore, widely used climate scenarios–such as the Shared Socioeconomic Pathways (SSP) developed by the IPCC (O Neill et al., 2017)–combine both physical and economic narratives, necessitating a comprehensive modelling framework to fully account for these interconnected aspects.
    In this paper, we introduce a holistic yet tractable “climate-dependent DFA” framework to support insurers and regulators in addressing climate-related challenges. Leveraging the interconnected structure of traditional DFA, our approach integrates both physical and economic dimensions of climate scenarios to capture impacts on assets and liabilities, which is particularly useful for areas like economic capital modelling, capital planning, and solvency monitoring. To ensure transparency and interpretability, we adopt tractable models, avoiding unnecessary complexity that could obscure key insights (Eling and Parnitzke, 2007).
    We also acknowledge that the impacts of climate change on individual risk types have been examined in various studies in the actuarial field. For example, Bourget et al. (2024) analyse the effect of climate change on pluvial flood occurrences. Outside the general insurance domain, Min et al. (2025) and Guibert et al. (2025) investigate climate change’s implications for mortality risk. Our paper extends those studies by providing a holistic modelling framework that assesses the financial impact of climate change on general insurers across their income statements, integrating multiple risk types and capturing the dependencies between them.
  • 2.
    Long-term nature of climate change: Climate change is inherently long-term, and it is usually challenging to distinguish the effects of global warming from internal variability over short horizons; indeed, internal variability ranges from inter-annual (e.g., El Niño–Southern Oscillation) to inter-decadal (e.g., Atlantic Multi-decadal Variability) (Chapter 11 of IPCC, 2021a, pp. 1546–1600). In addition, new regulations – such as IFRS S2 (IFRS, 2023) and AASB S2 (AASB, 2024) – require insurers to disclose the long-term financial impacts of climate change. A longer-term perspective is also essential for decisions by insurers and policymakers that extend across multiple years, including relocation planning (typically 8 to 40 years Bower and Weerasinghe, 2021) and reinsurance planning (influenced by cycles of about 9 to 11 years Meier and Outreville, 2006, 2010).
    An initial attempt to measure climate change effects on both assets and liabilities of general insurers simultaneously is presented in Gatzert and Özdil (2024), albeit limited to a one-year horizon. Our proposed framework utilizes a multi-year perspective within a DFA structure to assess the cumulative effects of climate change on general insurers. Unlike most solvency frameworks, where the required assessment is restricted to a one-year time horizon (e.g., Chapter 2, Section 2 of Delegated Regulation (EU) 2015/35 (EIOPA, 2015), Chapter 8 of SS17/16 (PRA, 2016)), a multi-year framework is essential for evaluating climate change impacts, given their inherently long-term nature.
  • 3.
    Uncertainty in climate change impacts: The future climate pathway is strongly shaped by societal developments, making scenario analysis a common tool for capturing this uncertainty (O Neill et al., 2017). While traditional DFA models use stochastic simulations to explore a range of outcomes (Kaufmann et al., 2001), they typically exclude key qualitative societal factors – such as environmental awareness, education, and political stability – that are critical for projecting climate change (O Neill et al., 2017). Furthermore, many widely used climate scenarios remain deterministic (O Neill et al., 2017; Bertram et al., 2020), whereas general insurance applications like economic capital modelling require both central and distributional projections.
    By enabling simulation of a distribution of financial outcomes across diverse climate pathways under the DFA setting, the framework supports assessments of both expected returns and risks, which are key considerations for insurers, actuaries, and regulators. While our analysis incorporates the economic and environmental assumptions underlying each scenario, evaluating the validity of those assumptions lies beyond the scope of this paper.
Our proposed framework is tailored to the general insurance sector by incorporating its unique features: it captures asset–liability interdependence through an interconnected structure and models the high variability of liability cash flows via DFA’s stochastic simulations. It also includes catastrophe reinsurance programs and accounts for the sensitivity of reinsurance premiums to capital constraints, which are also key factors affecting insurer profitability and solvency (see, e.g., Meier and Outreville, 2006). This differs from studies such as Min et al. (2025) and Guibert et al. (2025), which analyse climate change impacts on mortality risk, a risk type more pertinent to life insurance.
Finally, to demonstrate its application, we implement the framework in the Australian context. Australia’s geography and climate make it highly exposed to hazards such as bushfires, floods, and tropical cyclones, with risks expected to intensify under high-emission scenarios (Chapter 12 of IPCC, 2021b, pp. 1805–1812). These pressures may challenge insurers’ ability to underwrite and remain solvent. We calibrate the framework using real Australian data on insurance losses, macroeconomic indicators, and financial markets to provide realistic long-term insights. While this study focuses on Australia, the framework is adaptable to other markets, which we leave for future research.
1.3 Scope and implications of the paper
The primary objective of the methodology presented in this paper is to provide insights into industry-wide trends over a long-term horizon under the impact of climate change, rather than to inform short-term, insurer-specific decision-making. To this end, we develop a macro-level framework focused on national-level projections for the general insurance sector. Subnational modelling of hazards or socio-economic variables, and their use in granular decisions like portfolio steering or risk pricing, is beyond this paper’s scope. The goal is a generalised, scalable framework rather than one tailored to any insurer. However, individual insurers may extend it by adding model granularity and integrating tools such as proprietary catastrophe models (Mitchell-Wallace et al., 2017).
It is also worth mentioning that this paper provides an illustrative example of climate change impacts on the general insurance market to demonstrate the proposed framework’s application. However, it is not intended as a complete predictive analysis for any specific country (e.g., Australia). The aim is to show what is possible when integrating climate considerations into DFA and to highlight high-level insights into potential impacts on the general insurance sector.
Finally, this paper aims to provide a baseline model to assess the financial impact of climate change on the general insurance market, excluding the influence of short-term policy changes or regulatory interventions. However, relevant stakeholders can explore the effects of various government or regulatory actions by applying modifications to the proposed model and comparing the outcomes against the baseline.
1.4 Outline of the paper
In Section 2, we introduce the modelling framework for the proposed climate-dependent DFA. Section 2.1 then provides an overview of the structure of this framework. The design of its component modules is discussed from Section 2.2 through Section 2.4.
In Section 3, we present and analyse the numerical simulation outcomes – calibrated using Australian data – generated by the proposed framework. Section 5 concludes the paper, and Section 4 discusses the limitations and potential directions for future research.
2 Modelling framework for climate-dependent DFA: A DFA framework underpinned by climate inputs
In this section, we begin by providing an overview of our proposed framework’s structure in Section 2.1, starting with an outline of the main climate scenarios we borrowed, along with a discussion of their limitations. We then present the design of the component modules within the climate-dependent DFA framework, which collectively enable users to capture the comprehensive impacts of climate change on general insurers. In particular, Section 2.2 introduces the climate and hazard modules, which simulate climate variables along with the frequency and severity of natural catastrophes over the projection period. These outputs serve as climate inputs to the DFA. This is one of the key contributions of our framework, distinguishing it from traditional DFA models that typically generate future simulations without accounting for climate dynamics under different scenarios. Subsequently, Sections 2.3 and 2.4 describe the assets and liabilities modules, respectively, which project future investment returns and underwriting results under each climate scenario, based on outputs from the climate and hazards modules. Finally, Section 2.5 introduces the surplus module, which combines outputs from both the assets and liabilities modules, and presents key measures of general insurance financial performance.
The framework is designed to be general, with most component models described in Sections 2.2 to 2.5 being transferable to other insurance markets. However, while the overall structure of the hazard module is adaptable, the specific hazard types and climate covariates must be tailored to the region under study. In Section 2.2, we illustrate the modelling of individual hazard types and the selection of relevant climate covariates by using Australia as an example, while offering generalised guidance where applicable.
Figure 2.1: Modelling framework of climate-dependent DFA. The major hazard events depicted in the hazard module are for illustrative purposes only (based on the Australian context). For application in other settings, users may replace these with the predominant hazard types relevant to their selected country.
2.1 Model overview
The proposed climate-dependent DFA framework is illustrated in Figure 2.1. A key input to this framework is the set of climate scenarios, for which we adopt the Shared Socioeconomic Pathways (SSPs). These SSPs form a widely adopted framework in climate research and they are central to the IPCC’s climate risk assessments (O Neill et al., 2017). It should be noted, however, that the SSPs represent only one of several possible scenario frameworks. In particular, it is acknowledged that the SSPs may be succeeded by the Representative Emission Pathways (REPs) in the IPCC’s Seventh Assessment Report (Meinshausen et al., 2024), which aim to provide more policy-relevant, flexible, and up-to-date scenario representations. The integration of such alternative scenarios is left for future research.
Each SSP scenario is associated with a narrative, from which the economic growth rate at the technological frontier is derived (Dellink et al., 2017a). Starting from historical values, country-specific GDP projections are generated under the assumption that individual economies gradually converge toward this frontier. The convergence speed is determined by the degree of trade openness, as inferred from the scenario narratives (Dellink et al., 2017a). The emissions pathways consistent with the economic and environmental assumptions underlying each scenario are then used as inputs to climate models to produce projections of future climate at a much finer spatial resolution, typically at the level of gridded cells (Eyring et al., 2016).
The narratives of the selected representative climate scenarios are outlined below (O Neill et al., 2017):
  • 1.
    SSP 2.6 (“Sustainability”): Envisions a world characterised by progressive economic development and improving environmental conditions. The combination of low physical risk and sustainable economic growth results in low challenges for both mitigation and adaptation.
  • 2.
    SSP 4.5 (“Middle of the Road”)  [^1] We follow the naming convention for the SSP 4.5 scenario as used in the literature (O Neill et al., 2017). However, it should be noted that projections under SSP 4.5 might underestimate both emissions and temperature increases under current policies. For instance, while the best estimate of global temperature increase under SSP 4.5 is about 1.5 degree by 2030, this level of warming had already been reached by 2023 (Goessling et al., 2025).: Represents a development pathway aligned roughly with typical historical trends observed over the past century, leading to moderate mitigation and adaptation challenges.
  • 3.
    SSP 7.0 (“Regional rivalry”): Describes a world characterised by slowing economic growth and environmental degradation due to regional rivalries. Here, the combination of weak economic growth and elevated physical risk gives rise to high mitigation and adaptation challenges.
  • 4.
    SSP 8.5 (“Taking the highway”): Describes a world with rapid economic growth driven by competitive markets and innovation. Heavy reliance on fossil fuels, however, contributes to high physical risk and consequently high mitigation challenges, though strong economic growth leads to relatively low adaptation challenges.
A distinct feature of our framework is its interconnected design, allowing it to capture interactions among the various financial dimensions of general insurers. Beginning at the top of the diagram, we use the future climate and socioeconomic projections from each SSP scenario as inputs for modelling other variables within the DFA framework, following the cascading structure illustrated in the figure.
The projections of climate variables underlying each SSP scenario are used to estimate the frequency and severity of major natural hazard events in a selected country, thereby generating the market-level catastrophe insurance losses. On the liabilities side, these catastrophe loss estimates translate into insurance claims liabilities for general insurers, taking into account their reinsurance structures and market shares. On the assets side, socioeconomic projections under each SSP scenario, along with climate damage estimates from the hazards module, inform the simulation of investment returns. Finally, combining the resulting asset and liability forecasts allows us to derive the surplus of general insurers, representing an overall measure of market-wide financial performance.
In essence, the holistic nature of our proposed framework captures both direct and indirect climate change impacts. By recognising dependencies among various financial components, it offers a comprehensive view of climate change’s impact on the general insurance sector.
Remark 2.1.
Note that caution is warranted when interpreting the results derived from these scenarios, given the limitations of their underlying assumptions, particularly under high-emission pathways such as SSP 8.5. This scenario assumes continued economic growth without accounting for the risk of economic collapse (Keen et al., 2021). Moreover, labour productivity is treated as exogenous in SSP scenarios, overlooking potential negative impacts of climate change (Keen et al., 2021). In addition, recent studies suggest that the increasing duration and intensity of extreme heat under high-emission scenarios tend to reduce life expectancy (Guibert et al., 2025; Min et al., 2025), which could, in turn, negatively affect labour force growth. However, this potential impact is likewise not incorporated into the labour force growth assumptions in the current SSP frameworks. While SSP 8.5 also assumes strong investments in health, education, and highly engineered infrastructure (O Neill et al., 2017), which may mitigate climate-related productivity and labour force losses, the extent of such mitigation remains uncertain. In addition, the SSP scenarios do not capture abrupt or disorderly transition pathways and are therefore less suitable for stress tests involving severe transition shocks to the economy. Assessing or refining these assumptions lies beyond the scope of this paper and is left for future research.
Remark 2.2.
This paper focuses on the direct financial impacts of climate change on general insurers, as outlined in the modelling framework presented in Figure 2.1. While indirect effects, such as shifts in customer preference toward “greener” insurers or the potential rise in liability risks (e.g., lawsuits against commercial policyholders for environmental damage Alien, 2003; Bullock, 2022) that represent transition risks on the liability side are not included here, these are important areas for future research. In addition, we do not consider unidentified risks stemming from unforeseen responses of environmental and social systems to climate change, as these currently remain entirely unknown and unquantified (Rising et al., 2022). As quantitative and qualitative methodologies for assessing such risks continue to evolve, their integration into DFA frameworks may become more feasible.
2.2 Climate inputs
To incorporate the physical risk narratives embedded in various climate scenarios, we adapt climate projections from global climate models (subject to necessary modifications) to simulate the future frequency and severity of catastrophe events via the proposed climate and hazard modules. These outputs are then translated into the financial impacts on general insurers through assets and liabilities modules in the DFA framework based on its cascading structure. This approach serves as a major addition to traditional DFA models (Kaufmann et al., 2001), which typically rely on stationary hazard loss distributions derived from historical data. By integrating both historical information and scenario-based climate outlooks, our framework aims to produce forward-looking simulations of catastrophe insurance losses under the influence of climate change.
2.2.1 Climate module
The raw forecasts of climate variables (e.g., temperature, precipitation, and sea-level pressure) are derived from outputs of the Coupled Model Intercomparison Project Phase 6 (CMIP6). The findings based on the CMIP6 models, which play a crucial role in informing the IPCC Sixth Assessment Report (Lee et al., 2023). CMIP6 comprises a set of global climate model experiments that simulate historical, present and future climate conditions under IPCC’s SSP scenarios (Eyring et al., 2016). CMIP6 model outputs are typically provided as gridded datasets, representing climate variables across latitude–longitude grids over time, with spatial resolutions ranging from $1^{\circ}$ to $2.5^{\circ}$. These outputs are aggregated by averaging across grid cells within defined regions, with the selection of regions for each hazard type discussed in detail in Section 2.2.2.
One limitation of the raw outputs from CMIP6 models is that they are deterministic in nature. As highlighted in Section 1.2, it is essential for actuarial applications, especially capital modelling, to incorporate the stochastic variability (i.e., aleatoric uncertainty) of climate forecasts. Additionally, model outputs can exhibit biases relative to observations (often due to resolution discrepancies) (Haug et al., 2011; Maraun, 2013). Furthermore, uncertainties can also arise from limitations in the climate models used (i.e., model uncertainty) (Liu and Raftery, 2021).
To address these biases and both aleatoric and model uncertainty, we adopt the following procedure for simulating future climate variables, building on the approach of Liu and Raftery (2021):
  1. 1.
    Model uncertainty: We use an ensemble of CMIP6 models to capture differences among future climate forecasts (Liu and Raftery, 2021). This ensemble approach acknowledges that distinct models can yield varying projections.
  2. 2.
    Bias correction: For each ensemble member, we correct bias by comparing model backcasts to historical observations via the quantile mapping method, a simple but effective technique frequently used in the literature (Haug et al., 2011; Maraun, 2013; Sanabria et al., 2022). Specifically, using a quantile mapping approach with a linear transformation function (Piani et al., 2010; Qian and Chang, 2021), we estimate
    $\hat{\theta}_{q}\;=\;\hat{\beta}_{0}^{(m)}\;+\;\hat{\beta}_{1}^{(m)}\,\hat{\theta}_{q}^{(m)},$ (2.1)
    where $\hat{\theta}_{q}$ and $\hat{\theta}_{q}^{(m)}$ are the $q^{\text{th}}$ quantiles of the historical observations and model $m$ backcasts, respectively, over the same reference period.
  3. 3.
[…]
  4. 4.
    Future projections and simulations: For each simulation path in the future projection period, we randomly select a CMIP6 model $m$ to generate a deterministic forecast $\hat{\theta}_{t}^{(m)}$. We apply the bias correction as $\hat{\theta}_{t}^{(m),*}\;=\;\hat{\beta}_{0}^{(m)}+\hat{\beta}_{1}^{(m)}\,\hat{\theta}_{t}^{(m)}$, and then draw one trajectory of residuals $\tilde{z}^{(m)}_{t}$ to account for aleatoric uncertainty. The final simulated climate variable is thus:
    $\tilde{\theta}_{t}=\hat{\theta}_{t}^{(m),*}\;+\;\tilde{z}^{(m)}_{t}$. (2.2)
A schematic illustration of the process described above is shown in Figure 2.2.
Figure 2.2: An illustrative diagram of climate variable simulations
2.2.2 Hazards module
Based on the projected climate variables from the previous module, this section forecasts the frequency and severity of natural hazards. This constitutes a critical component of the DFA model, as the resulting hazard forecasts will be employed to model the general insurance assets and liabilities in subsequent sections. Numerous approaches exist for hazard modelling in the literature; however, as discussed in Section 1.2, balancing model interpretability and comprehensiveness is essential.
At one extreme, traditional Collective Risk Models (CRMs) (Klugman et al., 2012) offer a simplistic, intuitive means of modelling aggregate insurance losses, and it is also often used in traditional DFA applications (Kaufmann et al., 2001). Yet, their static assumption regarding insurance loss distributions neglects the dynamics introduced by climate change. At the other extreme, CAT models are sophisticated models that are usually capable of capturing the complex environmental process affected by climate change to generate hazard events based on advanced physical and mathematical models (Mitchell-Wallace et al., 2017). However, these proprietary models usually have complex structures with modelling details usually not accessible by general insurers, making them less comprehensible for insurers (Weinkle and Pielke Jr, 2017), leading to challenges in interpretability.
In light of the above considerations, we have opted for the weather-dependent CRMs (Haug et al., 2011) for modelling insurance losses. These models combine the high interpretability of traditional CRMs with the capability to incorporate climate effects by integrating meteorological variables in the modelling of insurance loss frequency and severity. In essence, the aggregate catastrophe loss ($\tilde{X}_{t}$) is modelled as:
$\tilde{X}_{t}=\sum_{i=1}^{I}\sum_{m=1}^{M_{t}^{(i)}}\tilde{X}_{t}^{(i),m},$ (2.3)
where $M_{t}^{(i)}$ is the number of event of hazard type $i$ in year $t$, and $\tilde{X}_{t}^{(i),m}$ is the insurance loss associated with the $m^{th}$ event of hazard type $i$ in year $t$. We further assume:
$M_{t}^{(i)}\sim\text{Poi}(\lambda(\bm{\Theta}_{t}^{(i)}));\;\;X_{t}^{(i),m}\sim\text{LN}(\mu(\bm{\Theta}_{t}^{(i)}),\sigma^{2}),$ (2.4)
where $\bm{\Theta}^{(i)}_{t}$ is a set of weather covariates for hazard type $i$, and $X_{t}^{(i),m}$ is the normalised catastrophe loss adjusted for both inflation and wealth exposure  [^2] The per-event normalised loss is defined as: $X_{t}^{(i),m}=\tilde{X}_{t}^{(i),m}\cdot\frac{\text{CPI}_{s}}{\text{CPI}_{t}}\cdot\frac{\text{Real GDP}_{s}}{\text{Real GDP}_{t}}$, which adjusts the nominal loss ($\tilde{X}_{t}^{(i)}$) for both the price level and total wealth in reference year $s$. Although our normalization technique is relatively simple, it aligns with conventional approaches in the literature (see, e.g., Pielke and Landsea, 1998; Vranes and Pielke Jr, 2009). While more granular factors could be incorporated into normalization (see, e.g., Crompton and McAneney, 2008; Pielke, 2021), our focus on future projections rather than historical trends – and on national-level rather than granular losses – supports the adoption of a more parsimonious approach.. The Poisson distribution aligns with common practice in modelling the frequency of hazard events (see, e.g., Jagger et al., 2008, 2011). Similarly, Log-Normal distribution is frequently used for modelling catastrophe losses due to its heavy-tailed nature (Kaufmann et al., 2001; McNeil et al., 2015), and it is also selected from a class of heavy-tailed candidates based on our data. The detailed discussion of the selected distributions based on the Australian data is provided in Online Appendix A.3.1. However, alternative distributions may provide a better fit for other datasets; hence, it is recommended to validate the distributional assumptions proposed here when applying the model to new data.
We further specify:
$\log(\lambda^{(i)}_{t})=\bm{\beta}_{(i)}^{\prime}\bm{\Theta}_{t}^{(i)};\;\;\mu^{(i)}_{t}=\bm{\alpha}_{(i)}^{\prime}\bm{\Theta}_{t}^{(i)},$ (2.5)
where the set of coefficients $\bm{\beta}_{(i)}$ and $\bm{\alpha}_{(i)}$ are estimated via regression. We do not explicitly adjust the scale parameter of the severity distribution in this study. However, it could be modelled as a function of climate covariates using GAMLSS regression (Stasinopoulos and Rigby, 2008). This is left for future research when stronger statistical and climate science evidence becomes available on the influence of climate covariates on the scale parameter.
We select the weather covariates based on the physical mechanisms driving each hazard type $i$ and validate them statistically. Here, we focus on major Australian hazards: flood, bushfire, tropical cyclones, storms, hailstorm, and East Coast Lows (Insurance Council of Australia, 2024). Below, we outline the candidate weather covariates for each hazard type, drawn from relevant literature on the associated physical processes. Although the data primarily cover Australia, the selection of candidate climate variables can be generalised to other regions, except for East Coast Lows, which is a unique hazard type occurring near the eastern coast of Australia.
The hazard model presented in this section is designed to capture industry-level trends consistent with the scope of this paper. For more granular decision-making at the level of individual insurers (e.g., portfolio management), a higher-resolution modelling framework may be required. A discussion of how the current hazard model can be extended to regional projections and integrated with CAT model outputs is provided in Online Appendix E. Nonetheless, incorporating such extensions should be carefully balanced against the trade-off between precision and interpretability depending on the intended business applications, as discussed in Section 1.2.
Bushfire
Bushfire risk depends on temperature, relative humidity, drought conditions, and wind speed (Sharples et al., 2016; Dowdy, 2018; Quilcaille et al., 2022). Generally, fire danger increases with higher temperatures, lower humidity, stronger winds, and more severe drought. The Fire Weather Index (FWI), derived from these factors, is commonly used to assess bushfire risk (Dowdy, 2018; Quilcaille et al., 2022). In particular, bushfire occurrence responds most strongly to extremes of the FWI rather than average conditions (Quilcaille et al., 2022). Hence, we choose:
$\bm{\Theta}_{t}^{\text{(BF)}}=\{\text{fwixx}_{t},\text{fwixd}_{t}\},$
where $\text{fwixx}_{t}$ is the annual maxima of fire weather index, and $\text{fwixd}_{t}$ is the number of days with extreme fire weather, which are two crucial statistics capturing FWI extremes (Quilcaille et al., 2022). The FWI data is subsequently averaged over the land surface area of the selected country to derive a national-level indicator.
Flood
Precipitation, particularly extreme precipitation, is a key driver of pluvial and river floods (Kodra et al., 2020). Accordingly, we choose:
$\bm{\Theta}_{t}^{\text{(FL)}}=\{R^{\text{x5}}_{t},R^{\text{x1}}_{t},R_{t}\},$
where $R_{t}$ represents the annual total precipitation, $R^{\text{x1}}_{t}$ is the largest one-day precipitation (usually denoted as rx1day), and $R^{\text{x5}}_{t}$ is the largest five-day cumulative precipitation (usually denoted as rx5day). Both $R^{\text{x1}}_{t}$ and $R^{\text{x5}}_{t}$ are commonly used as proxies for extreme precipitation in the IPCC report (Chapter 11 of IPCC, 2021a, pp. 1557–1561). Similarly, the precipitation data is then averaged across the land surface area of the selected country.
Tropical Cyclone and storms
The critical climate drivers for the formation and intensity of cyclones include sea surface temperature (SST) and sea level pressure (Meiler et al., 2022). Warm ocean water is an essential condition for storm formation; additionally, low sea level pressure contributes to storm development by causing warm, moist air to rise (Bureau of Meteorology, ). Therefore, storms typically form under conditions of high sea surface temperature and low sea level pressure. When the wind speed exceeds 119 km/h, the storm is classified as a cyclone (Bureau of Meteorology, ).
To capture the influence of sea surface temperature and sea level pressure on tropical cyclones and storms, we choose:
$\bm{\Theta}_{m}^{\text{(TC)}}=\{\overline{\text{SST}}_{m},\overline{\text{MSLP}}_{m}\},$
where $\overline{\text{SST}}_{m}$ denotes the monthly average sea surface temperature over the tropical cyclone basin near the selected country (e.g., the Australian Tropical Cyclone Basin for Australian applications)  [^3] Note: The cyclone basin refers to the area of tropical cyclone formation., and $\overline{\text{MSLP}}_{m}$ is the monthly average mean sea level pressure over the same region. The use of monthly averages for sea surface temperature and mean sea level pressure is also consistent with common practices in the literature, such as in the CAT model STORM (Bloemendaal et al., 2020).
Hailstorm
Atmospheric and near-surface temperatures are key drivers of hailstorm formation and intensity (Raupach et al., 2021). Rising near-surface temperatures increase low-level moisture and convective instability, potentially boosting hail frequency (Allen et al., 2014; Raupach et al., 2021). Higher atmospheric temperatures contribute to greater water vapor, intensifying hailstorms, but also raise the melting level, which can reduce hail frequency by melting smaller hailstones (Raupach et al., 2021). Thus, the overall impact varies locally. For instance, an increase in hail frequency is projected in Australia (Leslie et al., 2008; Allen et al., 2014; Raupach et al., 2021).
Accordingly, we choose:
$\bm{\Theta}_{m}^{\text{(H)}}=\{\bar{T}_{m}^{\text{NS}},\bar{T}_{m}^{\text{MT}}\},$
where $\bar{T}_{m}^{\text{NS}}$ is the monthly average near-surface temperature, and $\bar{T}_{m}^{\text{MT}}$ is the monthly average mid-tropospheric temperature.
East Coast Lows
An East Coast Low (ECL) is a type of mid-latitude cyclone that forms near the east coast of Australia, commonly referred to as the ECL identification region (Pepler et al., 2016a, b; Speer et al., 2021). However, the formation mechanisms of ECLs differ from those of tropical cyclones. While tropical cyclones develop over warm ocean waters, ECL formation is primarily driven by sea-surface temperature gradients (Pepler et al., 2016a). The ECL module presented here is specifically designed for Australian applications; for use in other countries, this module may require modification or removal.
Based on the above mechanism, we choose:
$\bm{\Theta}_{m}^{\text{(ECL)}}=\{\Delta\overline{\text{SST}}_{m},\overline{\text{SST}}_{m}\},$
where $\Delta\overline{\mathrm{SST}}_{m}$ represents the SST gradient near the east coast of Australia, defined as the average SST anomaly in the region $24^{\circ}$$41^{\circ}\mathrm{S}$ and $148^{\circ}$$155^{\circ}\mathrm{E}$ relative to the SST at the same latitudes and time between $160^{\circ}$$165^{\circ}\mathrm{E}$ (Pepler et al., 2016a).
Remark 2.3.
The hazard loss modelling in this paper does not explicitly incorporate potential government interventions. As discussed in Section 1.3, the primary aim of this work is to develop a general framework, rather than a complete predictive analysis for any specific country. As a baseline model, it can serve as a foundation for future studies to explore the potential impacts of various policy interventions.
Remark 2.4.
Forecasts of hazard-related losses are typically derived from historically calibrated relationships between climate variables and observed insurance losses. However, these relationships may change, especially under the impact of tipping points (Neal et al., 2025). Future research could improve hazard modelling by incorporating tipping point effects and conducting sensitivity analyses to account for the high degree of uncertainty in their timing, triggers, and impact magnitude (Lenton et al., 2008; Nordhaus, 2013).
Remark 2.5.
While our climate input modules account for scenario uncertainty, model uncertainty, and aleatoric uncertainty as discussed in previous sections, they do not incorporate parameter uncertainty and model inadequacy (i.e., the inherent limitations of models in fully representing real-world systems) as highlighted in Rising et al. (2022). These two sources of uncertainty are left for future research.
2.3 Assets and macro-economic variables
Beyond the modelling of hazard losses, DFA frameworks also commonly generate simulations of future macroeconomic variables and asset returns to capture the effects of changing economic and financial environments on the financial performance and position of general insurers (Coutts and Devitt, 1989; Kaufmann et al., 2001). To incorporate the influence of climate change on these factors, as discussed in Section 1.1, we introduce our assets and macro-economic variables modules. These modules aim to account for both the physical risks and the broader economic dimensions of climate change under different scenarios. This is achieved by extending traditional modelling approaches, drawing on relevant literature to reflect the long-term impacts of climate change on financial markets and economic conditions.
2.3.1 Inflation rates
General inflation rates can influence both the liabilities and assets of general insurers by affecting claims inflation and nominal interest rates (Kaufmann et al., 2001). Baseline inflation rates are modelled following the common approaches in literature by using the mean-reverting AR(1) process (see, e.g., Chen et al., 2021; Bégin, 2022):
$i_{t}=\mu_{i}+a_{i}(i_{t-1}-\mu_{i})+\sigma_{i}\epsilon_{i,t},$ (2.6)
where $i_{t}$ denotes the inflation rate at time $t$, $\mu_{i}$ is the long-run mean inflation, $a_{i}$ is the autoregressive parameter, $\sigma_{i}$ is the volatility, and $\epsilon_{i,t}$ represents a standard error term.
Studies have shown that historical fluctuations in weather conditions – such as temperature shocks and increased temperature variability – can exert inflationary pressures on food, energy, and service prices (Faccia et al., 2021; Mukherjee and Ouattara, 2021; Ciccarelli et al., 2023). This inflationary effect ultimately contributes to general inflation. Since climate change is expected to exacerbate weather fluctuations, it is crucial to account for its impact in modelling inflation rates (Kotz et al., 2024). To incorporate the influence of climate on inflation, we apply a climate overlay to the baseline inflation rates, following the methodology proposed by Kotz et al. (2024). To the best of our knowledge, the study by Kotz et al. (2024) is the first to quantitatively assess and project the effects of future climate change on both food and general inflation. Specifically, the climate-adjusted inflation rate is given by:
$i_{t}^{\text{Clim}}=i_{t}+i_{t}^{\text{Clim-Impact}},$ (2.7)
where $i_{t}^{\text{Clim-Impact}}$ captures the additional inflationary effects from climate change. Following the approach in Kotz et al. (2024), the monthly climate impact on inflation is modelled as:
$i_{m}^{\text{Clim-Impact}}=\sum_{L=0}^{11}(\alpha_{1+L}\Delta\bar{T}_{m-L}^{\text{NS}}+\beta_{1+L}\bar{T}_{m-L}^{\text{NS}}\cdot\Delta\bar{T}_{m-L}^{\text{NS}}),$ (2.8)
where $\bar{T}_{m}^{\text{NS}}$ denotes the monthly average near-surface temperature over the selected country and $\Delta\bar{T}_{m}^{\text{NS}}$ represents the deviation of future monthly averages from the 1990–2021 baseline. This formulation assumes a one-year lag effect. The term $\beta_{1+L},\bar{T}_{m-L}^{\text{NS}}\cdot\Delta\bar{T}_{m-L}^{\text{NS}}$ is introduced to capture the interaction effect whereby higher temperatures during hotter months lead to larger inflationary impacts (Faccia et al., 2021; Kotz et al., 2024). For future projections, the monthly average near-surface temperature will be sourced from the outputs of the climate module described in Section 2.2.1. The annual climate impact on inflation rates for year $t$ is then obtained by summing the monthly impacts (Kotz et al., 2024): $i_{t}^{\text{Clim-Impact}}=\sum_{m\in t}i_{m}^{\text{Clim-Impact}}$.
2.3.2 Risk-free interest rates
Drawing inspiration from Laubach and Williams (2003) and Holston et al. (2017), we model the real risk-free short-term interest rate as:
$r_{t}=\beta_{0}+\beta_{1}g_{t}+z_{t},$ (2.9)
which is closely related to the Ramsey’s equation (Ramsey, 1928), given by $r^{*}=\rho+\gamma g$, where $g$ denotes the growth rate of potential output, $\rho$ represents the rate of time preference, and $r^{*}$ denotes the natural rate of interest. A positive relationship between $r^{*}$ and $g$ is expected, as higher potential growth enhances future income prospects, reducing households’ incentives to save today and thereby placing upward pressure on natural rate of interest (Mongelli et al., 2022).
In our model, $g_{t}$ denotes the growth rate of potential real GDP. For calibration, these growth rates will be obtained from the World Bank Potential Growth Database (Kilic Celik et al., 2023). For future projections, $g_{t}$ will be derived from the GDP forecasts under each SSP scenario provided in the SSP database (Riahi et al., 2017). The residual term $z_{t}$ is assumed to follow an AR(1) process:
$z_{t}=\mu_{r}+\phi_{r}(z_{t-1}-\mu_{r})+\epsilon_{r}(t),$ (2.10)
which captures residual factors not explained by the growth rate. The nominal risk-free rate is then derived by incorporating inflationary effects using Fisher’s equation: $\tilde{r}_{t}=r_{t}+i_{t}^{\text{Clim}}$.
In summary, the key inputs for this model are the real GDP growth rate $g_{t}$ and the climate-adjusted inflation rate $i_{t}^{\text{Clim}}$ (as output from the inflation model; see Section 2.3.1). These inputs yield the nominal risk-free rate, $\tilde{r}_{t}$, as the final output.
Remark 2.6.
Choice of $g_{t}$: To mitigate the potential endogeneity issue, here $g_{t}$ is chosen as the growth rate of potential (full-capacity) GDP in the historical calibration; for future forecasts, $g_{t}$ will be derived from the potential real GDP forecasts underlying each SSP scenario (Dellink et al., 2017b). Therefore, the impact of monetary policy (through manipulation of $i_{t}$) on $g_{t}$ is limited, as it mainly affects short-term output gaps.
2.3.3 Equity module
We begin by considering the benchmark equity return model proposed by Ahlgrim et al. (2005), which is given by:
$r_{t}^{(S)}=\tilde{r}_{t}+x_{t},$ (2.11)
where $\tilde{r}_{t}$ is the nominal risk-free rates, and $x_{t}$ is the excess equity return.
Under traditional DFA or ESG frameworks (see, e.g., Wilkie, 1995; Ahlgrim et al., 2005; Chen et al., 2021; Bégin, 2022), excess equity returns are often modelled as independent stochastic processes, which is a self-contained approach that can enhance reliability in light of the considerable uncertainty surrounding exogenous variables over long-term horizons (Wilkie, 1995). However, relying solely on historical data limits the capacity to capture the forward-looking climate impacts and the evolving socio-economic conditions under different scenarios.
By contrast, factor models leverage a wide range of climate proxies – often at a granular level – to assess their influence on equity returns (see, e.g., Bansal et al., 2019; Hong et al., 2019; Görgen et al., 2020; Venturini, 2022). While these models are effective for empirical, in-sample analyses of individual or portfolio assets, their extensive data requirements and focus on asset-specific rather than market-level returns pose challenges for long-term projections, whereas ESG or DFA typically focuses on market-level returns.
To strike a balance between these two approaches, we propose a partial self-contained framework that incorporates forward-looking climate considerations without relying on overly granular external data. Inspired by the climate-economic literature (see, e.g., Karydas and Xepapadeas, 2022; Barnett, 2023), our method channels climate’s influence on equity returns through climate-damaged consumption. A flow diagram illustrating the simulation of excess equity returns is presented in Figure 2.3.
Figure 2.3: Illustrative diagram showing the simulation flow for equity excess returns
In the first step, we obtain the simulated nominal market insurance catastrophe loss $\tilde{X_{t}}$ from the hazard module (see Section 2.2.2). Next, these insurance losses are scaled to represent uninsured economic damage, yielding $\eta\tilde{X}_{t}$, where $\eta$ is the multiplier. The production available for consumption after climate damage is then computed as:
$C_{t}=Y_{t}-\eta\tilde{X}_{t},$ (2.12)
where $Y_{t}$ is the real GDP projection under each SSP scenario.
To transmit the effects of economic growth and climate damage to equity returns, excess total equity returns are modelled as a function of corporate earnings (based on operating profits) growth, $\Delta\text{OP}_{t}$, which is itself modelled as a function of consumption growth, $\Delta C_{t}$. Specifically, we have:
$\Delta\text{OP}_{t}=\alpha_{0}+\alpha_{1}\Delta C_{t}+\epsilon_{t}^{O},\;\;\;x_{t}=\beta_{0}+\beta_{1}\Delta\text{OP}_{t}+\epsilon_{t}^{\text{x}}$ (2.13)
where the parameters are estimated through regression using historical data. The random shocks to operating profit growth and excess equity returns are represented by $\epsilon_{t}^{\text{x}}\sim N(0,\sigma^{2}_{\text{x}})$ and $\epsilon_{t}^{O}\sim N(0,\sigma^{2}_{O})$, with their volatilities calibrated from historical observations.
For the brown sector, we further apply a transition stress overlay factor (Grippa and Mann, 2020) on its operating profit growth:
$\Delta\text{OP}_{t}^{B}=\Delta\text{OP}_{t}+\beta\Delta Y_{t}^{B},$ (2.14)
where $\Delta Y_{t}^{B}$ represents the change in brown energy production, and $\beta$ is the sensitivity of brown firms’ corporate profits to these changes. The adjusted operating profit growth for the brown portfolio is then used to simulate its excess equity returns based on (2.13).
Based on the outputs from the interest rate (Section 2.3.2) and equity modules, investment returns are calculated as: $r_{t}^{I}=w_{f}\tilde{r}_{t}+(1-w_{f})r_{t}^{(S)}$, where $w_{f}$ is the proportion of the portfolio allocated to risk-free assets.
In summary, the key inputs to this model are the nominal risk-free rate $\tilde{r}_{t}$, aggregate catastrophe losses $\tilde{X}_{t}$, real GDP projections, and brown energy production for each SSP scenario ($Y_{t}$ and $Y_{t}^{B}$), with $\tilde{r}_{t}$ and $\tilde{X}_{t}$ obtained from the respective interest rate and hazards modules. The model outputs are the equity returns for the general portfolio, $r_{t}^{(S,G)}$, and for the brown portfolio, $r_{t}^{(S,B)}$. By preserving the simplicity of traditional methods while integrating forward-looking climate damage projections, this partial self-contained approach captures key trends in evolving climate and socio-economic conditions without using high-dimensional external factors.
Remark 2.7.
The equity model presented here focuses solely on domestic market investments. In practice, however, insurers often hold foreign asset exposures. This limitation is less concerning over longer horizons, as the SSP framework assumes convergence in global economic growth (Dellink et al., 2017a), and market return differentials are expected to narrow over time due to arbitrage. The asset model is also intentionally simplified to align with the overall framework.
Another key assumption in the asset module is that short-term government bonds are considered free of default risk. This approach aligns with the risk-free treatment commonly adopted in conventional DFA studies (see, e.g., Kaufmann et al., 2001; D’Arcy and Gorvett, 2004; Consigli et al., 2018) and is consistent (in the Australian context) with the Australian Prudential Regulation Authority’s (APRA) Prescribed Capital Amount (PCA) framework for Australian sovereign bonds (APRA, 2023a). However, this assumption may warrant reconsideration in light of potential climate-induced sovereign downgrades as climate-related damages intensify (Klusak et al., 2023). The extent and significance of climate impacts on sovereign risk, however, vary across the literature and depend on the specific countries examined (Cevik and Jalles, 2022; Mallucci, 2022; Klusak et al., 2023). Future research could incorporate default risk into the modelling of government bond returns within DFA frameworks under climate change scenarios, leveraging existing findings to assess the potential material impact on the financial performance of general insurers.
In addition, when deriving net production after accounting for climate damage, we do not explicitly incorporate the reconstruction effects of natural disasters on GDP. The direction and magnitude of these effects remain largely inconclusive in the empirical literature. Albala-Bertrand (1993) found that, for most countries in their sample, short-term GDP growth increased following natural disasters, primarily due to the replacement of destroyed capital with more efficient ones, while the long-term growth rate remained unaffected. Skidmore and Toya (2002) reported that climatic disasters can also exert a positive influence on long-run economic growth in addition to short-term effects. In contrast, based on cyclone impacts, Hsiang and Jina (2014) rejected the hypothesis that disasters stimulate growth, showing that national income tends to decline and fails to recover in the short term. Similarly, Noy (2009) identified statistically significant short-term GDP losses, but with smaller magnitudes in developed economies due to their greater capacity to mobilize reconstruction resources. The theoretical framework of Hallegatte and Dumas (2009) further suggests that while disasters can affect the level of production, they do not alter the long-run growth rate, and the production response depends on the quality of reconstruction. Our assumption regarding economic damage shares some slight (and limited) similarity with Hallegatte and Dumas (2009) and Noy (2009), implying a short-run decline in production but no long-term effect on economic growth. However, the explicit incorporation and testing of alternative assumptions regarding the reconstruction effects on GDP are left for future research.
Finally, similar to other bottom-up approaches (e.g., Wasko et al., 2021), the derivation of economic damage in (2.12) accounts only for direct losses from acute climate risks. Future work could also integrate chronic risks into the climate-adjusted economic growth used in equity modelling, providing a more comprehensive representation of climate-related impacts.
2.4 Liabilities and premiums
Building on the outputs from the hazard and macroeconomic variable modules, the liabilities and premiums module introduced below aims to translate the impacts of climate change into the financial statements of general insurers.
2.4.1 Insurance costs
Drawing on the hazard module outputs (Section 2.2.2) and assuming an aggregate excess-of-loss reinsurance contract, the net catastrophe loss allocated to insurer $j$ is determined by:
$\tilde{X}_{t,(j)}^{\text{net}}=w_{j}\tilde{X}_{t}-\text{min}((w_{j}\tilde{X}_{t}-d_{(j)})_{+},L_{(j)}),$ (2.15)
where $\tilde{X}_{t}$ represents the gross CAT losses adjusted for CPI and GDP growth, $w_{j}$ denotes the market share of insurer $j$, and d(j)d{(j)} and $L_{(j)}$ are the inflation and GDP-adjusted reinsurance excess and limit levels for insurer $j$. The second term in (2.15) represents the recoverables from reinsurers based on the excess-of-loss contract.
In addition to catastrophe losses, another key component of DFA liability modelling is non-catastrophe losses (Kaufmann et al., 2001). We model non-catastrophe losses with a Tweedie distribution (Jørgensen and Paes De Souza, 1994), which is a commonly used distribution assumption for modelling non-catastrophe loss. Specifically,
$X_{t}^{\text{NC}}\sim\text{Tweedie}(\mu^{\text{NC}},\phi),$ (2.16)
where $\mu^{\text{NC}}$ is the location parameter and $\phi$ is the dispersion parameter, both calibrated using historical data. Using a Tweedie distribution implies that claim frequency follows a Poisson distribution, while claim severity follows a Gamma distribution, reflecting the typically high-frequency, low-severity nature of non-catastrophe losses. Some studies have investigated the influence of weather on non-catastrophe claims (e.g., McGuire, 2008; Haug et al., 2011; Scheel et al., 2013; Reig Torra et al., 2023), but these typically require high-resolution daily or monthly municipal-level data, exceeding the usual granularity of DFA. Moreover, the long-term effect of weather on non-catastrophe claims is uncertain. For instance, McGuire (2008) found that same-day precipitation increases motor claims frequency, whereas lagged precipitation decreases it possibly due to cleaner road conditions (Eisenberg, 2004; McGuire, 2008). Consequently, we do not explicitly incorporate climate impacts in our modelling of non-catastrophe losses, and the integration of climate covariates is left for future research as more data become available.
For projections, the aggregate non-catastrophe loss is computed as
$\tilde{X}_{t,(j)}^{\text{NC}}=X_{t}^{\text{NC}}\cdot\omega_{t}\cdot w_{j}\cdot\frac{\text{CPI}_{t}}{\text{CPI}_{s}},$ (2.17)
where $\omega_{t}$ is the total number of risks, $s$ denotes the reference year, and $\text{CPI}_{t}$ is the climate-adjusted CPI from the inflation module (see Section 2.3.1). The total number of risks is modeled as a linear function of population:
$\hat{\omega}_{t}=\hat{\omega}_{0}+\hat{\omega}_{1}\text{Pop}_{t},$ (2.18)
where $\text{Pop}_{t}$ is the projected population for each climate scenario, and the parameters $\hat{\omega}_{0}$ and $\hat{\omega}_{1}$ are calibrated on historical data via linear regression. Under these assumptions, climate change does not directly affect non-catastrophe losses; however, it still influences them indirectly through population growth and inflation.
Remark 2.8.
The exposure growth modelling presented here assumes that business volume grows in line with population growth and does not consider the potential loss of business volume due to premium increases driven by climate change. This simplification relies on the assumption that household income growth will generally keep pace with rising premiums. While this may be less concerning under scenarios such as SSP 8.5—where both economic growth and climate risk are high—or SSP 2.6—where climate risk is low—the issue of affordability may become more significant under scenarios with weak income growth but elevated climate risk (e.g., SSP 7.0). As a potential extension, one could incorporate the influence of GDP-per-capita growth relative to premium growth on premium affordability, and its implications on insurance demand.
In Australia, an initial effort to assess premium affordability in the context of climate change is being undertaken through the Insurance Climate Vulnerability Assessment project initiated by APRA (APRA, 2023b), with results expected by the end of 2025. Future research could build on these findings (or similar such findings in other jurisdictions) to incorporate the impact of premium affordability on business volume.
2.4.2 Insurance premiums
Based on the distribution assumptions of catastrophe and non-catastrophe losses, the insurance premium is then calculated using the standard deviations loading principle (Paudel et al., 2013, 2015; Tesselaar et al., 2020):
$\pi_{t,(j)}=\underbrace{\text{E}(\tilde{X}_{t,(j)})+\rho\sqrt{\text{Var}(\tilde{X}_{t,(j)})}}_{\text{CAT premium}}+\underbrace{\text{E}(\tilde{X}_{t,(j)}^{\text{NC}})+\rho\sqrt{\text{Var}(\tilde{X}_{t,(j)}^{\text{NC}})}}_{\text{Non-CAT premium}},$ (2.19)
where $\rho$ is the risk aversion parameter that reflects the level of insurer’s risk aversion towards the extreme nature of the risk (Paudel et al., 2013). The risk aversion parameter could be selected empirically. We adopt the assumed risk aversion parameter of 0.55 in Kunreuther et al. (2011) and Paudel et al. (2013), which is based on an empirical survey analysis conducted by Kunreuther and Michel-Kerjan (2011). The implication of this assumption on the projected premiums growth will also be examined in Section 3.2.3.
2.4.3 Reinsurance premiums
Based on the reinsurance structure specified in Section 2.4.1, the reinsurance premium is derived as:
$\pi_{t,(j)}^{RI}=\text{E}[\min((\tilde{X}_{t,(j)}-d_{(j)})_{+},L_{(j)})]+\rho\sqrt{\text{Var}(\min((\tilde{X}_{t,(j)}-d_{(j)})_{+},L_{(j)}))}.$ (2.20)
Similarly, the second term in (2.20) represents the surcharge on the premium above the expected value of the loss, which is dependent on the variability of the reinsurance losses.
Reinsurers typically cover the extreme tail of insurers’ risk portfolios through excess-of-loss coverage, making them particularly vulnerable to large natural catastrophes. Such events can strain reinsurance capital and trigger hard markets with higher premiums, which is a trend expected to intensify as climate change increases catastrophe losses (Tesselaar et al., 2020). Capital constraint theory offers a common explanation, suggesting firms prefer to accumulate surplus through higher premiums rather than raise costly external capital (Winter, 1988, 1994; Dicks and Garven, 2022).
To capture this relationship, we model reinsurance premiums as a function of reinsurance capital using a negative exponential form, inspired by Taylor (2008):
$\pi_{t,(j)}^{RI,*}=\max\left(\pi_{t,(j)}^{RI},\pi_{t,(j)}^{RI}e^{-k_{1}\cdot(S_{t-1}-S_{0})}\right),$ (2.21)
where $S_{t-1}$ denotes the solvency ratio at the end of period t1t-1, $S_{0}$ represents the reference (or steady-state) solvency ratio, and $k_{1}$ is the premium-to-solvency sensitivity parameter. The solvency ratio is defined as the ratio of reinsurance capital to premium (i.e., $S_{t-1}=K_{t-1}^{\text{Re}}/\pi_{t-1}^{RI,*}$), where the reinsurance capital is derived as:
$K_{t}^{\text{Re}}=(1+\tilde{r}_{t}^{(I)})\left(K_{t-1}^{\text{Re}}+\sum_{j=1}^{J}\pi_{t,(j)}^{RI,*}\right)-\sum_{j=1}^{J}\min\left[(\tilde{X}_{t,(j)}-d_{(j)})_{+},L_{(j)}\right].$ (2.22)
Here, we assume that reinsurers earn the same investment returns as direct insurers. The specification in (2.21) ensures that premiums increase as the solvency ratio decreases. Furthermore, due to the concave nature of the function, a decline in capital levels results in a more pronounced increase in premiums compared to the decrease in premiums when capital levels rise. This asymmetry aligns with the assumptions in Winter (1994), where insurers are assumed to be averse to the risk of bankruptcy.
Remark 2.9.
In modelling insurance premiums, we assume that insurance contracts are repriced annually to reflect changes in risk. While this assumption is appropriate for the scope of this paper, which is to provide a baseline framework, it should be noted that, in practice, premium adjustments may lag behind changes in risk due to underwriting cycles (see, e.g., Winter, 1994), market competition (see, e.g., Taylor, 2008), and rate regulation (see, e.g., Weiss et al., 2010; Taylor, 2008). Such pricing delays may further increase insolvency risk through periods of underpricing. Future research could incorporate these factors into the modelling of premium dynamics and assess their implications for the risk–return profile of general insurers under various SSP scenarios, relative to the baseline results presented in this paper.
We acknowledge that the reinsurance model presented here is a simplified representation, as it assumes a single reinsurer exclusively covering catastrophe losses in Australia and does not allow reinsurance prices to fall below their risk-based premium levels. In practice, the reinsurance market typically involves multiple participants, with reinsurers operating globally (BeInsure Media, 2025), offering both catastrophe and non-catastrophe coverage, and prices that can occasionally fall below their risk-based premium levels (Enz, 2002; Meier and Outreville, 2010). Despite its highly stylised nature, the model generally reflects the prudence that reinsurance prices are likely to rise above risk-based premiums as reinsurance capital tightens. However, it should not be interpreted as a comprehensive modelling of reinsurance market cycles due to its simplified assumption mentioned above. The potential impacts of climate change on the reinsurance cycle itself constitute an interesting topic for future research.
2.5 Surplus and performance measures
Based on the simulated quantities from the previous module, the market surplus process is derived as (Kaufmann et al., 2001):
$K_{t}=\sum_{j=1}^{J}K_{t}^{(j)}=\sum_{j=1}^{J}(1+\tilde{r}_{t}^{(I)})(K_{t-1}^{(j)}+\tilde{\pi}_{t}^{(j)})-(\tilde{X}_{t,(j)}^{\text{net}}+\tilde{X}_{t,(j)}^{\text{NC}}),$ (2.23)
where $K_{t}^{(j)}$ represents the surplus for entity $j$ at time $t$. Market insolvency is defined as $K_{t}<0$.
To determine the initial capital level $K_{0}^{(j)}$, the base capital requirement $\tilde{K}_{0}^{(j)}$ is calibrated to satisfy $\text{Pr}(K_{1}^{(j)}\leq 0)=0.5\%$, consistent with the solvency standards under Solvency II (Christiansen and Niemeyer, 2014). Recognizing that insurers typically maintain capital buffers above the minimum required capital to mitigate insolvency risk, we scale this base requirement using a target capital ratio $\tau$. Therefore, the final starting capital is thus computed as $K_{0}^{(j)}=\tau\tilde{K}_{0}^{(j)}$.
The financial performance of general insurers is typically evaluated using both returns and risk measures. For returns, we consider the median surplus $\text{med}(K_{t})$ – the median of the surplus distribution at time $t$ – alongside the expected surplus:
$\text{E}(K_{t})=\frac{1}{N}\sum_{n=1}^{N}K_{t}^{(n)},$ (2.24)
where $N$ is the total number of simulations. These two measures are commonly used returns metrics in DFA studies (see, e.g., Kaufmann et al., 2001).
For the risk measures, we use both insolvency probability and the deficit-given-insolvency ratio. The insolvency probability, a common DFA risk metric (Kaufmann et al., 2001), is calculated as the proportion of simulations yielding zero or negative capital:
$\text{P}(K_{t}<0)=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}(K_{t}^{(n)}\leq 0).$ (2.25)
The second measure we consider is the deficit-given-insolvency ratio, given by:
$\text{E}\left[\frac{-K_{t}}{L_{t}}\mid K_{t}<0\right],$ (2.26)
where $L_{t}=\sum_{j=1}^{J}\tilde{X}_{t,(j)}^{\text{net}}+\tilde{X}_{t,(j)}^{\text{NC}}$ represents the total claims liabilities at time $t$. This ratio measures the severity of the market deficit conditional on insolvency and is analogous to the Loss-Given-Default (LGD) metric used in reinsurance credit risk (Chen et al., 2020). Together, these risk measures provide a comprehensive assessment of both the likelihood and severity of adverse outcomes in the surplus process.
3 Application to the Australian general insurance market
After introducing the modelling framework for the climate-dependent DFA, this section presents numerical examples to demonstrate its application within the Australian general insurance market. Section 3.1 outlines the data sources and parameter calibrations used in the proposed framework. Section 3.2 analyses key simulation results from individual modules and discusses their potential financial implications for general insurers. Finally, Section 3.3 analyses and compares general insurers’ financial performance under different climate scenarios using both risk and return measures.
3.1 Data and calibration
3.1.1 Data sources
Climate data
The historical weather data used to calibrate the climate and hazard modules are sourced from ERA 5 reanalysis data (Copernicus Climate Change Service, 2024b), which combines past observations with current weather computer models to provide consistent estimates of atmospheric, land and oceanic climate variables from 1950 to the present. For future projections, we use the outputs from the CMIP6 models (Copernicus Climate Change Service, 2024a), which offers projected meteorological variables under various emissions scenarios up to 2100 at both monthly and daily resolutions. Additionally, the CMIP6 models generate backcasts of climate variables from 1850 to 2014, which –together with historical data – are used to calibrate the bias correction and aleatoric uncertainty models as detailed in Section 2.2.1. As our analysis focuses on the frequency and severity of hazards at the national level, we use the average values of climate observations across all gridded cells for calibration and projection purposes.
Hazards data
To calibrate the frequency and severity models for catastrophe insurance losses as detailed in Section 2.2.2, we draw on the ICA dataset (Insurance Council of Australia, 2024), which is maintained by the Insurance Council of Australia. This dataset covers all recorded natural disasters in Australia from 1967 to 2024, including variables such as disaster locations, start and end dates, and total insured damages. The insured damages data has been normalised for CPI and exposure prior to model calibration.
Macro-economic data
In Section 2.3.2, we calibrate the interest rate model using the RBA cash rates data (Reserve Bank of Australia, 2024) and potential GDP growth estimates from the World Bank Potential Growth Database (Kilic Celik et al., 2023). The RBA dataset covers cash rate targets and overnight cash rates from 1976 to 2023, while the World Bank Database offers annual Australian potential GDP growth from 1981 to 2021. Consequently, the calibration period spans 1981–2021.
In Section 2.3.1, we calibrate the inflation rates model with data from Australian Bureau of Statistics (2024b), which provides quarterly Consumer Price Index (CPI) information for Australia from 1948 to 2023. These data are aggregated to an annual scale to align with the DFA model’s time resolution.
For projections, we obtain socio-economic variables from the IIASA SSP database (Riahi et al., 2017), offering forecasts of GDP, population, and energy production at the five-year interval under multiple development paths at the country level. Spline interpolation (Erdogan, 2013) is then used to convert these projections to annual data.
Financial data
To calibrate the equity models described in Section 2.3.3, we employ Australian gross corporate operating profits and total returns data for the All-Ordinaries share index. The corporate operating profits dataset, sourced from Australian Bureau of Statistics (2024a), covers the total industry on a quarterly basis from 1994 to 2023. The All-Ordinaries index data, obtained from FactSet (2024), provides daily-to-annual total returns from 1992 to 2023.
Insurance market statistics
In addition to the data sources mentioned above, we draw on Australian insurance market statistics from the General Insurance Performance Statistics database (APRA, 2024c) and the General Insurance Institution-Level Statistics database (APRA, 2024a). These data inform our market assumptions and support the calibration of the non-catastrophe loss model (see Section 2.4.1). Because no direct non-catastrophe loss data are publicly available, we derive industry-level non-catastrophe losses by subtracting the ICA-recorded catastrophe losses from the total industry losses reported in the General Insurance Performance Statistics database (APRA, 2024a), after joining those two datasets.
The General Insurance Performance Statistics database contains quarterly aggregate financial data on Australian general insurers from 2002 to 2023, while the General Insurance Institution-Level Statistics database provides annual, institution-level financial information from 2005 to 2023. Both databases include key metrics such as insurance losses, premiums, equity bases, and the number of underwritten risks.
3.1.2 Calibration results
Due to the large number of parameters, the bias-correction and noise volatility parameters calibrated for selected climate variables in the climate module are presented in Online Appendix A.1 for illustrative purposes.
Table 3.1 presents the calibrated parameters for the selected model of each hazard type [^4] Each coefficient $\beta$ links a climate variable (superscript) to a distribution parameter (subscript). For example, $\beta_{(\lambda)}^{\text{rx5day}}$ denotes the sensitivity of frequency ($\lambda$) to rx5day, while $\beta_{(\mu)}^{\text{rx5day}}$ refers to its effect on the severity location parameter.. The full set of candidate models is detailed in Online Appendix A.3. Final models were chosen based on both physical relevance (see Section 2.2.2) and statistical performance. For example, in Table J, all flood frequency models show positive coefficients for precipitation, consistent with known flood mechanisms. Model 3, using rx5day as an extreme precipitation proxy, performed best by AIC and BIC and its covariate is statistically significant at the $5\%$ level. Thus, Model 3 is selected. As another case, although the atmospheric temperature coefficient is not statistically significant at the $5\%$ or $10\%$ levels, we retain it in the hailstorm model due to the theoretical importance of temperature effects (see Section 2.2.2).
Hazards Parameters Values Data (for calibration) Data (for projection)
Flood $\beta_{(\lambda)}^{\text{rx5day}}$ 0.037** ERA 5 re-analysis; ICA data CMIP 6
$\beta_{(\mu)}^{\text{rx5day}}$ 0.035* ERA 5 re-analysis; ICA data CMIP 6
Bushfire $\beta_{(\lambda)}^{\text{mfwixx}}$ 0.084** ERA 5 re-analysis; ICA data CMIP 6
Cyclones $\beta_{(\lambda)}^{\text{SST}}$ 1.213*** ERA 5 re-analysis; ICA data CMIP 6
Storms $\beta_{(\lambda)}^{\text{SST}}$ 0.348** ERA 5 re-analysis; ICA data CMIP 6
$\beta_{(\mu)}^{\text{SST}}$ 0.239 ERA 5 re-analysis; ICA data CMIP 6
East Coast Low $\beta_{(\lambda)}^{\Delta\text{SST}}$ 2.189* ERA 5 re-analysis; ICA data CMIP 6
Hails $\beta_{(\lambda)}^{T^{\text{NS}}}$ 0.211*** ERA 5 re-analysis; ICA data CMIP 6
$\beta_{(\lambda)}^{T^{\text{MT}}}$ -0.079 ERA 5 re-analysis; ICA data CMIP 6
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table 3.1: Calibration of key parameters: Hazard modules
The calibrated parameters for the real risk-free interest rate and inflation rate models are shown in Table 3.2. The positive impact of potential GDP growth on real rates (as captured by the parameter $\beta_{1}$) aligns with the expectations discussed in Section 2.3.2. For temperature impacts on inflation, the coefficients $\alpha_{1+L}$ and $\beta_{1+L}$ are taken from Kotz et al. (2024), who calibrated these parameters across 121 countries to evaluate climate-related effects on inflation.
Model Parameters Values Data (for calibration) Data (for projection)
Real rates $\beta_{1}$ ($\partial{r_{t}}/\partial{g_{t}}$) 2.206** RBA cash rate; SSP database (IIASA)
World Bank Potential Growth Database
$\mu_{r}$ -0.0005 same as above.
$\phi_{r}$ 0.478** same as above.
$\sigma_{r}$ 0.025 same as above.
Inflation $\alpha_{1+L}$, $\beta_{1+L}$ (Kotz et al., 2024) CMIP 6
$\mu_{i}$ 0.0517*** ABS CPI data
$a_{i}$ 0.713*** ABS CPI data
$\sigma_{i}$ 0.0309 ABS CPI data
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table 3.2: Calibration of key parameters: Macro-economic variables
Table 3.3 presents the calibrated equity model parameters. $\alpha_{1}$ (sensitivity of operating profit growth to consumption growth) and $\sigma_{O}$ (its standard deviation) are calibrated using Australian operating profit data (Australian Bureau of Statistics, 2024a) and consumption growth data (World Bank Group, 2024). $\beta_{1}$ (sensitivity of excess equity returns to operating profit growth) and $\sigma_{X}$ (standard deviation of excess equity returns) are calibrated using operating profit data (Australian Bureau of Statistics, 2024a) and All-Ordinaries total returns (FactSet, 2024). Results indicate positive relationships between consumption and profit growth, and between profit growth and equity returns, consistent with expectations as discussed in Online Appendix C.1.
To calibrate the sensitivity of operating profits in brown firms  [^5] Here, we define brown firms as those operating within the oil and gas sector. to planned changes in production under various climate scenarios (sourced from the SSP database (Riahi et al., 2017)), we adopt a simplified approach akin to that used in Grippa and Mann (2020). We estimate the impact of changes in output on operating profits for a representative firm in the Australian energy sector by considering its fixed and variable costs. These results are then extrapolated to other oil and gas firms within the sector. Woodside Energy Group Ltd (WDS) is selected as the representative firm, given its dominant market share of $66\%$ in the Australian energy sector as of March 2024 (FactSet, 2024). The impact of output changes is calculated as the average percentage change in operating profit per $1\%$ change in production, based on the historical financial statements available from 2014 to 2021 in FactSet. While this approach is intentionally simplified, general insurers may refine it to better align with their portfolio compositions. Enhancements could include integrating more granular transition risk metrics, potentially incorporating proprietary data sources.
Parameters Values Data (for calibration) Data (for projection)
$\alpha_{1}$ ($\partial{\Delta{OP}_{t}}/\partial{\Delta C_{t}}$) 3.824* World Development Indicators (World Bank); SSP database (IIASA)
ABS Business Indicators CMIP 6
$\sigma_{\text{O}}$ 0.083 same as above.
$\beta_{1}$ ($\partial{x_{t}}/\partial{\Delta{OP}_{t}}$) 0.047 ABS Business Indicators; All-Ordinaries index (Factset) SSP database (IIASA)
$\sigma_{\text{x}}$ 0.103 same as above.
$\beta^{\text{(B)}}$ ($\partial{\Delta{OP}_{t}^{B}}/\partial{\Delta Y_{t}^{B}}$) 1.768 Income statements of WDS (Factset) SSP database (IIASA)
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table 3.3: Calibration of key parameters: Equity returns
Table 3.4 presents the market assumptions underlying our simulations, derived from general insurance market statistics (APRA, 2024b) and financial disclosures from individual insurers. In the projections, both excess and limit levels are adjusted for changes in GDP and CPI across future periods and under different climate scenarios.
Size of Insurers Numbers Market Shares Excess (normalised) Limit (normalised)
Large 4 20% \$1000 million \$600 million
Medium 4 3% \$150 million \$90 million
Small 12 0.67% \$33 million \$20 million
Table 3.4: Market assumptions
Table 3.5 summarises the additional assumptions used in the simulations. The target capital ratio is determined as the general insurance industry’s average ratio of eligible equity to the Minimum Capital Requirement (MCR), based on data from the APRA General Insurance Institution-Level Statistics database (APRA, 2024a). The uninsured-to-insured loss ratio is used to scale catastrophe insurance losses from the hazard module to uninsured economic damage, serving as an input to the equity model. This ratio is calculated as the average proportion of uninsured to insured losses over the historical period from 1985 to 2023, using data from the EM-DAT database (EM–DAT, 2023), since the ICA dataset records only insurance losses. It should also be noted that this ratio represents only the proportion of uninsured to insured damages to assets, which are primarily properties and motor vehicles that tend to have relatively high insurance coverage due to regulatory and mortgage requirements (Melser et al., 2024), during natural disasters. It does not, however, reflect the overall insurance penetration of the broader economy. Moreover, this ratio may vary across countries and can be affected by differences in the quality and reporting protocols of hazard impact data across regions (EM-DAT, 2023).
Parameters Type Values Source
$\rho$ Risk-aversion 0.55 Kunreuther et al. (2011); Paudel et al. (2013)
$\tau$ Target capital ratio 1.75 APRA (2024a)
$\eta$ Uninsured-to-insured loss ratio 1.22 EM–DAT (2023)
$w_{rf}$ Allocation to risk-free assets $0\% 66$6The assumption is that the majority of bonds are invested in risk-free government securities, as insurers typically prefer government bonds over corporate bonds (OECD, 2023). Additionally, cash and deposits are considered nearly risk-free assets OECD (2023)
$w_{B}$ Allocation to brown assets $3\%$ Gatzert and Özdil (2024)
Table 3.5: Other assumptions
Remark 3.1.
The data used to calibrate the hazard models are normalised to 2022 level for both CPI and exposure. For future projections, we adjust hazard losses using projected CPI and real GDP growth under each climate scenario. This differs from the normalisation technique in McAneney et al. (2019) (also adopted by ICA), which adjusts for dwelling numbers, values, sizes, and building standards based on historical data. Therefore, our approach assumes dwelling values across regions grow in line with nominal GDP, with no changes in dwelling size or building standards under different climate scenarios. We acknowledge this is a limitation, as climate change may affect dwelling numbers and values unevenly across regions, and adaptation measures could alter dwelling size and standards. These aspects are left for future research.
3.2 Key simulation results from individual modules
3.2.1 Climate and hazards
The simulated climate variables derived from the CMIP6 model outputs (Copernicus Climate Change Service, 2024a) are presented in Appendix A. Most variables show an upward trend, especially under high-emission scenarios, indicating a general increase in climate risk. The projections also exhibit notable inter-annual variability and a high degree of uncertainty, both of important drivers for the results shown in the downstream modules.
Figure 3.1: Simulation results of normalised hazard losses. Solid lines represent the average simulation paths under different climate scenarios, while dashed lines denote the $5^{\text{th}}$ and $95^{\text{th}}$ percentiles. Results are derived from simulated climate variables and calibrated hazard models (see Sections 2.2.2 and 3.1.2). The simulations reveal an increasing trend in both the mean and volatility of hazard losses for most hazard types under high-emission scenarios.
Based on the simulated climate variables, catastrophe losses for each hazard type are generated using the calibrated relationships between climate variables and hazard frequency and severity as described in Section 3.1.2. Overall, the projected normalised losses exhibit an increasing trend for most hazards, particularly under high-emission scenarios, with notable increases in both the mean and the upper tails of the loss distributions. In addition, the projected losses exhibit substantial interannual variability. For example, a pronounced peak in simulated flood losses is observed around 2047, particularly at the $95^{\text{th}}$ percentile. This peak is primarily driven by the corresponding increase in extreme precipitation (see Online Appendix B.1), which may be explained by the projected rise in the frequency of extreme ENSO events under high-emission scenarios reported in the literature (see, e.g., Cai et al., 2015; Heede and Fedorov, 2023). Moreover, the results reveal considerable uncertainty, driven by both the variability in simulation of climate variables as shown in Appendix A and the heavy-tailed nature of catastrophe losses. The rising mean and volatility are expected to place upward pressure on both insurance and reinsurance premiums, as well as shocks to capital over time.
Remark 3.2.
Despite the simplified structure of our hazard models, the simulation results generally align with physical risk trends reported in the literature. Projected increases in bushfire (with the upward trend more clearly illustrated in the extended simulation results presented in Online Appendix B.2) and flood losses are consistent with previous findings (Chapter 12 of IPCC, 2021b, pp. 1805–1811), reflecting intensified extreme fire weather and increased extreme precipitation across much of Australia.
For tropical cyclones, existing literature suggests a decline in total cyclone numbers but a rise in high-intensity events (Chapter 12 of IPCC, 2021b, pp. 1809–1810). As our frequency models are based on the ICA dataset, which only includes catastrophe-level hazards, the projected increase reflects trends in severe tropical cyclones and is broadly consistent with the literature.
Regarding hailstorms, a net increase in losses is projected. This is primarily driven by the influence of near-surface temperatures on storm frequency (see Table 3.1), and aligns with expected increases in hailstorm frequency (Leslie et al., 2008; Allen et al., 2014; Raupach et al., 2021).
For East Coast Lows, simulations show a slight decreasing trend, driven by changes in sea-surface temperature gradients. This result is consistent with studies projecting reduced East Coast Low frequency under high-emission scenarios (Pepler et al., 2016a, b; Speer et al., 2021).
Our focus is on broad trends in hazard losses rather than precise magnitudes, as the analysis aims to capture industry-wide patterns. For granular decision-making—such as capital allocation or portfolio steering—insurers can use proprietary claims data to tailor hazard models to their specific exposures.
3.2.2 Investment returns
The simulation results for the cumulative returns on the total investment portfolios, presented on a logarithmic scale, are shown in Figure 2(a)  [^7] Here, the log-scale cumulative investment return is defined as: $r^{\text{cum}}_{t}=\log\Bigl(\prod_{s=1}^{t}\bigl(1+r_{s}\bigr)\Bigr)$. This can also be written as: $r^{\text{cum}}_{t}=\log\Bigl(\prod_{s=1}^{t}\bigl(1+r_{s}\bigr)\Bigr)=\log\Bigl(\frac{V_{0}\prod_{s=1}^{t}(1+r_{s})}{V_{0}}\Bigr)=\log\Bigl(\frac{V_{t}}{V_{0}}\Bigr)$, where $V_{t}$ is the total investment value at time $t$.. The uncertainty bounds in the figure account for historical fluctuations in interest rates and equity returns, as well as the variability in climate-related damage to production available for consumption, as illustrated in Figure 3(a) and derived via (2.12).
The observed trends in cumulative investment returns are primarily driven by the economic growth assumptions underlying each SSP scenario (Figure 2(b)). Under SSP 8.5, the highest cumulative investment return is projected, driven by its robust economic growth assumption (O Neill et al., 2017), followed by SSP 2.6. Conversely, SSP 7.0 exhibits the weakest cumulative investment return, influenced by both its slow economic growth assumptions and high catastrophe damages, in line with its narrative (O Neill et al., 2017). The relatively high investment returns under SSP 8.5 and SSP 2.6 are likely to accelerate capital accumulation. By contrast, the lower returns, coupled with high catastrophe losses in SSP 7.0 (see Figures 3(b) and 3.1), are expected to slow the pace of capital accumulation under this scenario. A sensitivity analysis using alternative economic damage assumptions has also been conducted (see Online Appendix G), indicates broadly consistent patterns with those reported in this section.
A breakdown analysis of investment returns for the general and brown portfolios is provided in Online Appendix C.2. Overall, the transition stress remains relatively mild, as the SSP scenarios considered do not assume abrupt or disorderly transitions, as noted in Remark 2.1.
Remark 3.3.
The projected damage ratios from our climate-dependent DFA model are also compared with estimates from existing literature; a detailed discussion can be found in Online Appendix D and G.
Figure 3.2: Simulated (log) compounded investment returns (a) and projected compounded real GDP growth in Australia (b). Panel (a) shows the simulated compounded returns on the total investment portfolios generated from the economic growth assumptions and the simulated hazard losses underlying each scenario. Panel (b) presents compounded real GDP growth projections derived from the SSP database (Riahi et al., 2017).
Figure 3.3: Ratios of uninsured (a) and total (b) catastrophe losses to GDP. The results are obtained by dividing the scaled outputs from the hazard module (see Section 2.3.3) by the corresponding GDP projections under each climate scenario.
3.2.3 Premiums and underwriting losses
The normalized gross premiums associated with catastrophe (CAT) coverage are presented in Figure 4(a). An overall increasing trend is found, driven by the rising risk of most hazards as shown in Figure 3.1. Gross CAT premiums reach their highest levels under SSP 8.5, particularly in the later projection period, followed by SSP 7.0, SSP 4.5, and SSP 2.6, in line with the physical risk narratives underlying those scenarios as described in Section 2.1. The increase in premiums may also create affordability issues and, as noted in Remark 2.8, affect insurance demand. This effect is noted as a limitation of the current paper but is encouraged for future research. In addition to this general upward trend, significant inter-annual variability is observed, reflecting the internal climate variability of the underlying climate variables as shown in AppendixA.
Moreover, the proportion of CAT premiums relative to total general insurance premiums is expected to increase relative to the historical levels, with a more pronounced increase under high-emission scenarios. This is illustrated in Figure 4(b), which compares projected CAT premium proportions with historically observed CAT loss proportions (used here as a proxy for historical CAT premium proportions) derived from the General Insurance Performance Database (APRA, 2024b) and the ICA dataset. These findings suggest that catastrophe losses will have an increasingly significant impact on the underwriting performance of general insurers under the influence of climate change. Consequently, managing CAT exposures will become a more critical component of portfolio management for general insurers.
A similar pattern to the gross insurance premiums emerges in the reinsurance premiums (Figure 5(a)), which show a general upward trend. The highest projected premium occurs under SSP 8.5, followed by SSP 7.0, while SSP 2.6 is expected to have the lowest premium.
To illustrate the role of reinsurance capital constraints, Figure 5(b) presents the average relative difference between the solvency-sensitive reinsurance premiums (in (2.20)) and the base premiums (in (2.21)). The largest uplift appears under SSP 7.0, likely due to low investment returns coupled with high insurance losses (Sections 3.2.1 and 3.2.2), resulting in slower capital accumulation and prolonged tight capacity. In contrast, SSP 8.5 exhibits the smallest uplift despite experiencing the highest catastrophe (CAT) losses, possibly because its stronger investment returns (see Section 3.2.2) accelerate capital accumulation and shorten tight-capacity periods. A similar pattern is also observed in the primary general insurance market (see Section 3.3).
The results also suggest that the percentage growth in reinsurance premiums exceeds that of primary insurance, reflecting both rising physical risks and adjustments to tightening reinsurance capital. The modelled uplift relative to the base level is considered as conservative due to the simplified assumptions outlined in Remark 2.9. Nevertheless, it is assumed that primary insurers will continue to purchase reinsurance despite higher prices; however, rising premiums may dampen demand, although this effect could be partly offset by the strengthening risk-transfer demand as climate risk intensifies (Lloyd’s, 2025). The dynamics of these effects are left for future research.
Figure 3.4: Gross premiums (normalised) associated with catastrophe cover, shown as dollar values (a) and as a proportion of total gross premiums (b). Panel (a) presents the normalised values of calculated gross catastrophe (CAT) premiums based on (2.19). Panel (b) shows the projected CAT premiums as a proportion of total premiums, compared with historically observed CAT losses as a share of total insurance losses, based on data from the APRA database and the ICA dataset. Both panels indicate an increasing trend in CAT premiums–both in dollar terms and as a share of total premiums–particularly under high-emission scenarios.
Using the calculated premiums and simulated catastrophe losses from the hazard module (Section 3.2.1), we derive the simulated underwriting losses, with normalized results presented in Figure 3.6. On average, the simulated underwriting loss remains below zero (indicating a positive underwriting profit) and is relatively consistent across scenarios, which could be explained by the premium loadings. At higher quantiles, however, underwriting losses remain similar across scenarios initially but rise substantially in later periods under high-emission pathways (SSP 8.5 and SSP 7.0) due to escalating climate risks. This upward shift in the tail of underwriting losses can be attributed to the increasing volatility and extreme percentile of hazard losses observed under high-emission scenarios (see Section 3.2.1). Consequently, while downside liability impacts on financial performance appear moderate at first, they are expected to intensify over longer projection horizons in the high-emission scenarios.
Figure 3.5: Total reinsurance premiums (normalised) (a) and relative difference between total and base reinsurance premiums (b). Panel (a) shows the normalised values of calculated total reinsurance premiums (including mark-up), averaged across simulation paths, based on (2.21). The results exhibit trends similar to those observed for gross CAT premiums. Panel (b) presents the relative difference between total reinsurance premiums and base reinsurance premiums, with the latter calculated using (2.20). The largest premium uplift is projected under the SSP 7.0 scenario.
Figure 3.6: Simulations of underwriting losses (normalised). Underwriting losses are calculated by subtracting the net premium received (i.e., gross premium received minus reinsurance premiums paid) from the net simulated insurance losses (i.e., gross insurance losses net of reinsurance recoverables). Solid lines represent the average simulation paths across climate scenarios, while dashed lines indicate the $5^{\text{th}}$ and $95^{\text{th}}$ percentiles. Across climate scenarios, the results are similar at the mean level but diverge at higher quantiles, with greater losses projected under high-emission scenarios.
3.3 Risk and returns measures
Finally, the risk and return measures for the general insurance market can be derived from the surplus, which is calculated using outputs from the individual modules (see (2.23)). Figures 7(a) and 7(b) show the expected and median surplus, which are our return measures, across different SSP scenarios. Under SSP 8.5, surplus is highest, followed by SSP 2.6, whereas SSP 7.0 yields the lowest surplus.
To investigate the drivers of these surplus trends, we compare the expected and median surplus with cumulative investment returns at their average and median paths, revealing that differences in cumulative investment returns largely explain the observed surplus patterns (see Appendix B). These findings align with earlier results showing that underwriting profits or losses are similar at mean levels across scenarios (Section 3.2.3), making investment growth the primary driver of mean and median surplus trends. Since investment returns are predominantly affected by the economic growth assumptions underlying each scenario (Section 3.2.2), the divergent economic growth paths in different SSP scenarios are expected to be the key drivers of expected returns to the general insurance industry.
Figure 8(a) shows market insolvency probabilities under various scenarios, a common risk metric in DFA studies (Kaufmann et al., 2001). SSP 7.0 exhibits the highest insolvency rates, followed by SSP 4.5 and SSP 2.6. Although SSP 7.0 does not incur the highest hazard losses, its poor investment returns (Section 3.2.2), substantial reinsurance premium increases (Section 3.2.3), and relatively high catastrophe (CAT) losses (Section 3.2.1) collectively erode profits and constrain capital accumulation, leading to high insolvency probabilities once both physical and economic aspects of the climate change are considered. These outcomes align with the physical and economic narratives underlying the SSP 7.0 scenario (Section 2.1).
Figure 3.7: Expected (a) and median (b) market surplus (log scale) derived from surplus simulations. Both the expected and median surplus trajectories reflect the economic growth assumptions underlying each climate scenario.
Interestingly, while SSP 8.5 is typically linked to the highest physical risks in climate science (Chapter 11 of IPCC, 2021a, pp. 1517–1605), it yields the lowest projected insolvency probabilities when both physical and economic factors are modelled. This outcome stems from strong investment returns outpacing underwriting losses early in the projection horizon (see Figures 2(a) and 3.6). As a result, capital accumulates rapidly, particularly in the early years. SSP 8.5 also shows the highest average Compound Annual Growth Rate (CAGR) of market surplus  [^8] The Compound Annual Growth Rate (CAGR) of the insurance surplus over the projection horizon $T_{h}$ is given by: $\text{CAGR}_{h}=\left(\frac{K_{T_{h}}}{K_{0}}\right)^{\frac{1}{T_{h}}}-1$. This measure, commonly used in finance to assess the average growth rate of an investment portfolio (Grimm, 2023), is employed here to evaluate the rate of capital accumulation across different projection horizons.(see Appendix C), which could help insurers absorb potential losses at later stages. These findings align with the SSP 8.5 narrative of “robust economic growth”, which leads to “low adaptation challenges” except in extreme case (O Neill et al., 2017).
However, when insolvency occurs, SSP 8.5 results in the most severe impacts in later projection horizons, as shown by the market deficit-given-insolvency ratios (see Figure 8(b)). These ratios, which indicate the share of unpaid claims, suggest policyholders tend to face the greatest losses under SSP 8.5 in tail events. This is likely due to higher underwriting losses in the distribution tails later in the projection (Section 3.2.3; Figure 3.6), despite similar early-horizon losses across scenarios. Additionally, given the limitations of SSP 8.5 (Section 2.1), insurers should remain cautious about unmodeled risks such as economic collapse from climate tipping points.
SSP 2.6 offers a more balanced risk-return profile, with the second-highest expected surplus, second-lowest insolvency risk, and modest market deficits in insolvency. It also features the lowest projected catastrophe-related premiums, potentially enhancing affordability and insurance coverage against climate risks. These results align with SSP 2.6’s narrative of “sustainable economic development” and “improving environmental conditions,” leading to “low mitigation and adaptation challenges” (O Neill et al., 2017).
Figure 3.8: Market insolvency probabilities (a) and market deficit-given-insolvency ratios  [^10] Note that the market deficit-given-insolvency ratio exhibits high volatility across all climate scenarios in the early projection horizon. This is likely due to the relatively small number of simulation paths that become insolvent at early stages (see Figure 8(a)), which leads to greater fluctuation in the calculated ratios.(b). Panel (a) shows the proportion of simulations in which total market capital becomes negative, while Panel (b) presents the ratio of market deficit to total claims liabilities, conditional on market insolvency. Solid lines represent the average simulation paths across climate scenarios, and dashed lines indicate the $5^{\text{th}}$ and $95^{\text{th}}$ percentiles. The SSP 8.5 scenario exhibits the lowest market insolvency probability but the highest market deficit-given-insolvency ratio.
Remark 3.4.
The results on surplus and insolvency probability presented in Section 3.3 are based on the assumption that no dividends are paid from the surplus, consistent with the paper’s objective of establishing a baseline model. We also conducted a sensitivity analysis to examine the impact of different dividend policies on surplus and insolvency probability. Overall, although the magnitudes of surplus and insolvency probability vary under different dividend policies, the scenario rankings with respect to surplus and insolvency probability under varying dividend payout assumptions remain consistent with the baseline findings reported in this section. It is also found that the SSP 8.5 scenario shows greater sensitivity to higher dividend payouts, as larger dividends erode the capital buffer to absorb high catastrophe losses under this scenario. Detailed results and discussions are provided in Online Appendix F.
4 Discussion
The modelling framework necessarily adopts simplifying assumptions to manage the complexity of climate risk. While this offers a practical starting point, it introduces several key limitations that warrant attention. It also presents several opportunities for future extension, as discussed below.
4.1 Reliances and limitations
The first limitation of our analysis is its reliance on economic and environmental assumptions in the SSP scenarios, which have inherent limitations (see Remark 2.1). In addition, post-disaster reconstruction effects have not been explicitly incorporated into adjusting the GDP assumptions for climate damage (see Remark 2.7). Validating and refining these assumptions, as well as assessing their impact on insurers, offers a promising research avenue that might require interdisciplinary collaboration.
Secondly, our hazard loss forecasts rely on historically calibrated relationships between climate variables and insurance losses, which may shift over time. Future research could adopt a more forward-looking approach to reflect these changes (see Remark 2.4).
Thirdly, the framework is limited to domestic disasters and excludes global spillover effects. However, climate-related economic damage can potentially spread via international trade (Neal et al., 2025). Modelling these linkages would offer a more complete view of risks to local insurers.
Additionally, as noted in Remark 2.8, the framework does not consider potential business volume loss from rising premiums. Also, as highlighted in Remark 3.1, the analysis does not incorporate possible changes in future building standards or dwelling sizes that may result from adaptation measures when projecting hazard losses. Incorporating these factors in future work would provide a more comprehensive assessment of the long-term financial impacts on general insurers.
Finally, as discussed in Remark 2.5, while the framework incorporates several sources of climate-related uncertainty, it does not capture uncertainty arising from parameters and model inadequacy (see Remark 2.5). The unidentified risks are also not quantified in our framework due to their unquantifiable and truly unknown nature (see Remark 2.2). Users should therefore remain mindful of these additional risks and uncertainties when interpreting the model outputs.
4.2 Extensions and future work
Our framework offers several avenues for extension. First, while this paper assesses climate risks at an industry-wide level, it provides a foundation for individual insurers to develop firm-specific, climate-dependent DFA models aligned with their unique risk profiles. In addition, the current modelling framework does not differentiate between lines of business. Although this simplification is appropriate for the study’s macro-level focus, future research could extend the framework to incorporate multiple business lines, which would be particularly useful for decision-making at the individual insurer level, such as capital allocation (Myers and Read Jr, 2001).
Second, as noted in Remark 2.3 and Remark 2.9, the model serves as a baseline that does not incorporate the effects of government interventions, market competition, or regulatory actions. While certain reinsurance market dynamics have been considered in this study, it should not be regarded as a comprehensive modelling of reinsurance cycles due to the simplified assumptions outlined in Remark 2.9. Future research could incorporate these factors and compare the outcomes with this baseline to assess their impact on the financial performance of general insurers under different climate scenarios.
Third, on the liability side, the model focuses on climate impacts on catastrophe losses. Future research could extend this by incorporating climate effects on non-catastrophe losses as relevant data become available.
Moreover, as noted in Remark 2.2, transition risks on the liability side are not yet included. Their integration could be explored in future work once quantification methods are more established in the literature.
Finally, the modelling approach is broad but simplified. Future studies may adopt more advanced component models for greater precision, such as the proprietary catastrophe (CAT) models. These enhancements, however, must be balanced against the trade-off between model complexity and parsimony, which is an essential consideration in DFA.
5 Conclusions
This paper makes two key contributions. First, we propose a climate-dependent DFA framework that integrates climate risk into traditional Dynamic Financial Analysis. While previous studies examine the impacts of climate change on assets and liabilities separately, a unified framework remains underexplored. General insurers commonly use DFA to assess overall financial performance, yet conventional models do not account for climate risk. Our framework addresses this gap by leveraging the interconnected structure of DFA to capture the complex interactions of climate impacts on both assets and liabilities. Additionally, it incorporates a multi-year perspective to reflect the long-term nature of climate change. The uncertainty surrounding climate risks is addressed through stochastic simulations within climate scenario analysis, enabling the assessment of both risk and return dimensions. Finally, by incorporating the unique characteristics of general insurers, our framework provides realistic insights into the financial consequences of climate change on the insurance market.
Second, we conduct an extensive empirical study to assess the long-term impact of climate change on the Australian general insurance market, illustrating the practical use of our framework. In climate science, SSP 8.5 carries the highest physical risk, followed by SSP 7.0, SSP 4.5, and SSP 2.6 (Chapter 11 of IPCC, 2021a, pp. 1517–1605), a ranking reflected in our hazard simulations and CAT premium projections. However, when both economic and physical factors are considered, SSP 7.0, which combines high catastrophe risk with poor economic growth, emerges as the most detrimental scenario for insurers, leading to low returns and high insolvency risk. In contrast, SSP 8.5 delivers the highest returns and lowest insolvency risk due to strong economic growth, but also results in the largest market deficits upon insolvency, increasing risks to policyholders.
These findings highlight the need for insurers to prepare for scenarios involving both high catastrophe risk and weak economic growth in strategic areas such as business planning, capital management, and reinsurance. Insurers should also account for the tail-end financial risks under high-emission scenarios like SSP 8.5. Regulators, in turn, should ensure adequate capital buffers and collaborate with governments on contingency plans, such as bail-out mechanisms, to manage systemic risks from potential insurer insolvencies. Further development of the proposed climate-dependent DFA framework might support these efforts.
Data and Code
The R code used in this study is available at https://github.com/agi-lab/climate-dependent-DFA, and the supporting datasets are hosted at https://zenodo.org/records/18521235. Instructions for data access and code execution are provided on the GitHub page.
Due to licensing restrictions, we cannot share the Total Returns series of the All-Ordinaries Shares Index; users should obtain it directly from FactSet. For those without access, pseudo data simulated from the calibrated equity model is provided for calibration.
Similarly, we cannot distribute Woodside Energy Limited’s 2014–2023 financial statements from FactSet. Users can retrieve them via FactSet or compile them manually from public reports at https://www.woodside.com/investors/reports-investor-briefings. For models using restricted datasets, the same parameters are applied in the code to ensure consistency with the results presented.
Acknowledgements
This research was supported under Australian Research Council’s Discovery Project (DP200101859) funding scheme. The views and opinions expressed in this paper are solely those of the authors and do not reflect those of their affiliated institutions.
The authors would like to extend their sincere gratitude to Sharanjit Paddam for his constructive feedback and insightful comments, which significantly improved the quality of this paper and provided valuable guidance for potential future research directions.
Earlier versions of this paper were presented at the 2024 Australasian Actuarial Education and Research Symposium (AAERS) in Melbourne (Australia), All Actuaries Summit 2025 in Sydney (Australia), the 2025 28th International Congress on Insurance: Mathematics and Economics held in Tartu (Estonia), as well as at research seminars hosted by the School of Risk and Actuarial Studies at UNSW. The authors are grateful for constructive comments received from colleagues who attended those events.
The authors also gratefully acknowledge the Actuaries Institute of Australia for awarding the Melville Prize and the Carol Dolan Prize to an earlier version of this paper, which was tailored for an industry audience and presented at the All Actuaries Summit 2025.
Declarations of interest
None
A Simulation results of climate variables
Figure A presents the simulated climate variables derived from the CMIP6 model outputs (Copernicus Climate Change Service, 2024a). The results suggest that most climate variables, except for SST gradients, exhibit an upward trend, particularly under high-emission scenarios, indicating an overall increase in climate risk.
In addition to this long-term trend, the average simulation path also shows notable inter-annual variability. This variability, likely driven by internal climate processes (e.g., El Niño cycles) captured by the CMIP6 models (Jain et al., 2023), may contribute to annual fluctuations in insurance premiums and costs as shown in the corresponding sections in the paper.
Furthermore, the projections display a high degree of uncertainty. This uncertainty, also highlighted in other studies (see, e.g., Wu et al., 2022), arises from both the climate model uncertainty and the aleatoric uncertainty incorporated in our climate module (see Section 2.2.1). Given that such uncertainty, especially at the upper percentiles, is crucial for actuarial applications and capital modelling, these findings underscore the importance of accounting for uncertainty in climate projections, as discussed in Section 1.2 and 2.2.1.
Figure A: Simulation results for climate variables. Results are generated using CMIP6 model outputs and the simulation methodology described in Section 2.2.1. Solid lines represent the average simulation paths under different climate scenarios, while dashed lines indicate the $5^{\text{th}}$ and $95^{\text{th}}$ percentiles. The simulations illustrate both long-term trends and inter-annual variability, as well as the uncertainty associated with the projections.
B Relationships between market surplus and investment returns
Figure A: Expected (a) and median (b) market surplus vs. average (a) and median (b) compounded investment returns (log scale). The results indicate a strong correlation between market surplus and compounded investment returns at both the mean and median levels.
C Mean Compound Annual Growth Rate (CAGR) of market surplus
Figure A: Mean Compound Annual Growth Rate (CAGR) of market surplus, by projection horizon and climate scenario. This metric reflects the rate of capital accumulation over different projection periods. The highest growth rate is observed under the SSP 8.5 scenario during the early projection horizon.
Online Appendix
This online appendix accompanies the paper “Dynamic Financial Analysis (DFA) of General Insurers under Climate Change” by Benjamin Avanzi, Yanfeng Li, Greg Taylor, and Bernard Wong.
Included here are supplementary details on model calibration results for the climate and hazard modules (see Section A); supplementary simulation outputs from the hazard module (see Section B); supplementary details on the equity model (see Section C); a benchmarking analysis of projected damage ratios (see Section D); discussion on potential extensions of the current model to multi-regional projections, portfolio management, and integration with CAT modelling outputs (see Section E); a sensitivity test on the impacts of dividend payout policies (see Section F); and a sensitivity test on the impacts of alternative economic damage assumptions on investment returns (see Section G).
A Model calibration results: supplementary details
A.1 Climate model supplementary calibration results
The results presented here show the calibrated bias-correction and noise volatility parameters for selected climate variables within the climate module. The detailed calibration methodology is provided in Section 2.2.1 of the main paper.
Model names $\hat{\beta}_{0}^{(m)}$ $\hat{\beta}_{1}^{(m)}$ $\hat{\sigma}_{(m)}$
ACCESS-CM2 2.257 0.883 1.271
CanESM5-CanOE 2.484 0.889 1.354
CESM2 -0.065 0.987 1.256
CMCC-CM2-SR5 -2.18 1.085 1.405
CNRM-CM6-1 0.735 0.974 1.321
CNRM-ESM2-1 0.868 0.942 1.186
FGOALS-f3-L 0.944 0.98 1.294
FGOALS-g3 -1.132 1.039 1.23
INM-CM4-8 1.345 0.971 1.199
INM-CM5-0 2.222 0.948 1.195
IPSL-CM6A-LR -0.139 1.055 1.428
MCM-UA-1-0 1.329 0.899 1.401
MIROC-ES2L -2.734 1.045 1.255
MIROC6 -0.85 0.931 1.325
MPI-ESM1-2-LR 1.441 0.916 1.288
MRI-ESM2-0 -0.046 0.96 1.375
NorESM2-$M$ -0.858 1.036 1.267
Table A: Near-surface temperature (monthly-average): Calibrated bias-correction coefficients ($\hat{\beta}_{0}^{(m)}$ and $\hat{\beta}_{1}^{(m)}$), and standard deviation of noise ($\hat{\sigma}_{(m)}$)
Model names $\hat{\beta}_{0}^{(m)}$ $\hat{\beta}_{1}^{(m)}$ $\hat{\sigma}_{(m)}$
ACCESS-CM2 0.926 1.045 1.622
CanESM5-CanOE -4.122 0.924 1.32
CESM2 -4.897 0.881 1.491
CNRM-CM6-1 6.783 1.04 1.412
CNRM-ESM2-1 5.617 1.025 1.366
INM-CM4-8 12.879 1.361 1.346
INM-CM5-0 14.56 1.384 1.502
IPSL-CM6A-LR 1.294 1.001 1.409
MCM-UA-1-0 -0.307 0.982 1.368
MIROC-ES2L -10.139 0.675 1.488
MIROC6 -4.526 0.86 1.369
MPI-ESM1-2-LR 1.847 1.036 1.475
MRI-ESM2-0 3.861 1.102 1.448
NorESM2-$M$ -6.233 0.838 1.408
UKESM1-0-$L$ 2.097 1.066 1.596
Table B: Air temperature (monthly-average): Calibrated bias-correction coefficients ($\hat{\beta}_{0}^{(m)}$ and $\hat{\beta}_{1}^{(m)}$), and standard deviation of noise ($\hat{\sigma}_{(m)}$)
Model names $\hat{\beta}_{0}^{(m)}$ $\hat{\beta}_{1}^{(m)}$ $\hat{\sigma}_{(m)}$
ACCESS-CM2 -2.418 0.996 0.316
CNRM-CM6-1 4.446 0.801 0.263
FGOALS-f3-L 2.266 0.924 0.35
FGOALS-g3 0.792 0.955 0.279
INM-CM4-8 1.414 0.944 0.239
INM-CM5-0 0.636 0.994 0.255
MCM-UA-1-0 2.26 0.881 0.378
MIROC-ES2L -5.092 1.204 0.356
UKESM1-0-$L$ 1.157 0.943 0.262
Table C: Sea-surface temperature (monthly-average): Calibrated bias-correction coefficients ($\hat{\beta}_{0}^{(m)}$ and $\hat{\beta}_{1}^{(m)}$), and standard deviation of noise ($\hat{\sigma}_{(m)}$)
Model names $\hat{\beta}_{0}^{(m)}$ $\hat{\beta}_{1}^{(m)}$ $\hat{\sigma}_{(m)}$
ACCESS-CM2 -17.133 1.306 8.675
ACCESS-ESM1-5 -4.806 1.239 8.346
CanESM5 -0.279 1.234 9.263
CMCC-CM2-SR5 57.288 1.052 7.644
CMCC-ESM2 53.705 0.959 10.141
EC-Earth3 -45.561 1.726 8.791
FGOALS-g3 8.191 1.323 8.697
GFDL-ESM4 5.471 1.261 8.534
INM-CM4-8 17.222 1.582 9.137
INM-CM5-0 -9.668 1.898 8.932
IPSL-CM6A-LR 42.053 1.074 8.376
KACE-1-0-G -35.045 1.209 8.592
MIROC6 30.162 0.911 9.741
MPI-ESM1-2-HR -42.055 1.476 8.781
MPI-ESM1-2-LR -23.581 1.359 8.883
MRI-ESM2-0 -20.99 1.108 8.844
NorESM2-$M$ 40.105 0.992 8.963
TaiESM1 50.755 1.001 9.536
Table D: Fire Weather Index: Calibrated bias-correction coefficients ($\hat{\beta}_{0}^{(m)}$ and $\hat{\beta}_{1}^{(m)}$), and standard deviation of noise ($\hat{\sigma}_{(m)}$)
Model names $\hat{\beta}_{0}^{(m)}$ $\hat{\beta}_{1}^{(m)}$ $\hat{\sigma}_{(m)}$
ACCESS-CM2 2471.323 0.975 146.064
CanESM5-CanOE 18345.06 0.819 136.658
CESM2 14433.794 0.857 154.145
CMCC-CM2-SR5 14248.639 0.859 144.904
CNRM-ESM2-1 2981.637 0.97 144.572
INM-CM4-8 -12552.876 1.125 133.775
INM-CM5-0 -25169.454 1.25 145.112
MCM-UA-1-0 37431.853 0.629 125.137
MIROC-ES2L 6250.218 0.942 179.028
MIROC6 624.259 0.998 169.48
MPI-ESM1-2-LR -8807.205 1.087 170.452
MRI-ESM2-0 -23938.494 1.234 185.582
NorESM2-$M$ 15219.903 0.85 155.486
UKESM1-0-$L$ 10695.377 0.894 142.076
Table E: Mean sea-level pressure (MSLP): Calibrated bias-correction coefficients ($\hat{\beta}_{0}^{(m)}$ and $\hat{\beta}_{1}^{(m)}$), and standard deviation of noise ($\hat{\sigma}_{(m)}$)
Model names $\hat{\beta}_{0}^{(m)}$ $\hat{\beta}_{1}^{(m)}$ $\hat{\sigma}_{(m)}$
access-cm2 2.811 1.156 18.276
access-esm1-5 4.2 1.051 17.659
bcc-csm2-mr -22.791 1.417 20.799
canesm5 -12.108 1.517 18.868
cmcc-esm2 6.167 1.007 19.607
ec-earth3 -12.368 1.2 19.394
gfdl-esm4 -14.868 1.321 16.222
inm-cm4-8 11.73 0.985 18.008
inm-cm5-0 3.857 1.115 18.983
ipsl-cm6a-lr -5.697 1.305 17.26
kace-1-0-g 16.459 0.953 16.631
miroc6 -45.993 1.728 19.952
mpi-esm1-2-hr 14.079 0.875 17.57
mpi-esm1-2-lr -3.133 1.185 17.843
mri-esm2-0 20.245 0.835 16.37
noresm2-lm -19.198 1.364 18.09
noresm2-$m$ -17.776 1.263 18.665
Table F: Largest five-day cumulative precipitation (rx5day): Calibrated bias-correction coefficients ($\hat{\beta}_{0}^{(m)}$ and $\hat{\beta}_{1}^{(m)}$), and standard deviation of noise ($\hat{\sigma}_{(m)}$)
Model names $\hat{\beta}_{0}^{(m)}$ $\hat{\beta}_{1}^{(m)}$ $\hat{\sigma}_{(m)}$
ACCESS-CM2 -0.199 1.31 0.498
CNRM-CM6-1 0.399 1.053 0.485
FGOALS-f3-L -1.589 0.722 0.513
FGOALS-g3 -1.713 1.047 0.529
INM-CM4-8 0.63 0.765 0.501
INM-CM5-0 0.074 0.71 0.481
MCM-UA-1-0 -0.229 1.044 0.47
MIROC-ES2L -0.264 0.774 0.437
UKESM1-0-$L$ -0.021 0.636 0.462
Table G: Sea-surface temperature gradients: Calibrated bias-correction coefficients ($\hat{\beta}_{0}^{(m)}$ and $\hat{\beta}_{1}^{(m)}$), and standard deviation of noise ($\hat{\sigma}_{(m)}$)
A.2 Normality tests of climate model residuals
The Shapiro–Wilk test is applied to assess the normality of the climate model residuals. The resulting p-values are reported below. For most ensemble members, the p-values exceed the 5% threshold, indicating insufficient statistical evidence to reject the null hypothesis of normality at the conventional 5% significance level.
Figure A: p-values from Shapiro-Wilk test of climate model residuals: Near-surface temperature (left), air temperature (middle), and sea-surface temperature (right)
Figure B: p-values from Shapiro-Wilk test of climate model residuals: Fire Weather Index (left) and extreme precipitation (rx5day) (right)
A.3 Hazard model supplementary calibration results
The results presented here show the full calibration outcomes for the hazard models, including all candidate models considered for each hazard type. The methodology for calibrating these parameters is provided in Section 2.2.2 of the main paper.
A.3.1 Selection of catastrophe loss distributions
Table H reports the AIC and BIC values for the candidate heavy-tailed severity distributions (Embrechts et al., 2013). The Log-normal distribution achieves the best fit according to both criteria.
Distributions AIC BIC Rank (AIC) Rank (BIC)
Log-Normal 9366.825 9366.825 1 1
Weibull 9405.851 9412.683 3 3
Pareto 9379.475 9386.307 2 2
Cauchy 9560.989 9567.821 4 4
Table H: Goodness-of-fit for severity distributions: AIC and BIC
As for the frequency of catastrophe events, we assess the suitability of the Poisson distribution assumption using a two-sided dispersion test (Cameron and Trivedi, 1990). Overdispersion or underdispersion would indicate that an alternative distribution (e.g., Negative Binomial) may be more appropriate. The estimated dispersion parameters across all hazard types are close to one, and the p-values show no statistically significant evidence of over- or under-dispersion at the 5% level. Hence, the Poisson distribution is deemed appropriate for modelling event frequency in the current application.
Hazard type Dispersion estimates z-statistic p-value
Flood 1.061 0.376 0.707
Bushfire 1.117 0.608 0.543
Tropical Cyclones 0.968 -1.693 0.090
ECL 0.990 -1.333 0.183
Hailstorms 1.144 1.112 0.266
Table I: Dispersion test results for Poisson frequency models across different hazard types.
A.3.2 Flood
Model 1 Model 2 Model 3* Model 4 Model 5 Model 6
(Intercept) 2.177-2.177** 4.295-4.295** 3.714-3.714** 4.326-4.326** 3.666-3.666** 4.191-4.191**
($0.004$) ($0.002$) ($0.001$) ($0.005$) ($0.002$) ($0.002$)
Precipitation 0.0040.004** $0.000$ $0.000$
($0.005$) ($0.967$) ($0.902$)
rx1day 0.0810.081** 0.0830.083+ $0.041$
($0.002$) ($0.098$) ($0.515$)
rx5day 0.0370.037** 0.0350.035+ $0.020$
($0.001$) ($0.099$) ($0.487$)
AIC $138.6$ $136.1$ $136.0$ $138.1$ $138.0$ $137.6$
BIC $142.7$ $140.1$ $140.0$ $144.1$ $144.1$ $143.7$
Log.Lik. 67.306-67.306 66.026-66.026 65.998-65.998 66.025-66.025 65.991-65.991 65.788-65.788
F $7.838$ $9.906$ $10.511$ $4.948$ $5.285$ $5.269$
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table J: Flood: candidate frequency models
Model 1 Model 2 Model 3* Model 4 Model 5 Model 6
Precipitation 0.0030.003+ $0.000$ 0.001-0.001
($0.086$) ($0.942$) ($0.826$)
rx1day 0.0710.071* $0.076$ $0.004$
($0.049$) ($0.335$) ($0.963$)
rx5day 0.0350.035* $0.040$ $0.034$
($0.029$) ($0.170$) ($0.339$)
AIC $2000.5$ $1999.6$ $1998.7$ $2001.6$ $2000.6$ $2000.6$
BIC $2006.1$ $2005.2$ $2004.3$ $2009.1$ $2008.1$ $2008.1$
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table K: Flood: candidate severity models
A.3.3 Bushfire
Here, mfwixx represents the annual maximum of the Fire Weather Index averaged across all gridded cells in Australia, while xfwixx denotes the highest annual maximum of the Fire Weather Index across all gridded cells. Similarly, mfwixd represents the average duration of extreme fire weather across all gridded cells, whereas xfwixd denotes the longest duration of extreme fire weather observed across all gridded cells in Australia.
Model 1* Model 2 Model 3 Model 4 Model 5 Model 6
mfwixx 0.0840.084** 0.0920.092**
($0.002$) ($0.001$)
xfwixx $0.004$ 0.005-0.005
($0.418$) ($0.482$)
mfwixd 0.0720.072** 0.1090.109**
($0.001$) ($0.007$)
xfwixd $0.012$ 0.014-0.014
($0.111$) ($0.262$)
AIC $118.3$ $127.9$ $119.7$ $120.1$ $126.3$ $120.8$
BIC $122.3$ $132.0$ $125.8$ $124.2$ $130.3$ $126.9$
Log.Lik. 57.130-57.130 61.968-61.968 56.858-56.858 58.053-58.053 61.137-61.137 57.394-57.394
F $9.889$ $0.657$ $5.433$ $10.195$ $2.547$ $5.470$
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table L: Bushfire: candidate frequency models
As all coefficients in the bushfire severity model are statistically insignificant and the stationary Log-Normal distribution yields the lowest AIC and BIC values, we adopt the stationary Log-Normal distribution to model bushfire severity.
Model 1 Model 2 Model 3 Model 4 Model 5 Model 6
mfwixx $0.016$ $0.008$
($0.652$) ($0.822$)
xfwixx $0.008$ $0.007$
($0.461$) ($0.532$)
mfwixd 0.003-0.003 $0.029$
($0.931$) ($0.603$)
xfwixd 0.004-0.004 0.012-0.012
($0.652$) ($0.495$)
Num.Obs. $37$ $37$ $37$ $37$ $37$ $37$
AIC $1527.0$ $1526.6$ $1528.6$ $1527.2$ $1527.0$ $1528.7$
BIC $1531.8$ $1531.5$ $1535.0$ $1532.0$ $1531.8$ $1535.1$
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table M: Bushfire: candidate severity models
A.3.4 Tropical cyclones
Model 1* Model 2 Model 3
Intercept 33.005-33.005*** 545.650545.650*** 367.462367.462***
(<<0.001) (<<0.001) (<<0.001)
SST 1.2131.213*** 0.9890.989***
(<<0.001) (<<0.001)
MSLP 0.005-0.005*** 0.004-0.004***
(<<0.001) (<<0.001)
AIC $257.3$ $258.0$ $244.1$
BIC $266.3$ $267.0$ $257.7$
Log.Lik. 126.625-126.625 127.005-127.005 119.067-119.067
F $27.719$ $36.066$ $16.122$
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table N: Tropical cyclones: candidate frequency models
Model 1 Model 2 Model 3 Model 4
SST 79.405-79.405
($0.189$)
$\text{SST}^{2}$ $1.560$
($0.194$)
MSLP 2.022-2.022
($0.342$)
$\text{MSLP}^{2}$ $0.000$
($0.343$)
cs(SST) 0.886-0.886 1.096-1.096*
($0.129$) ($0.048$)
cs(MSLP) 0.001-0.001
($0.346$)
AIC $1612.2$ $1613.0$ $1614.5$ $1613.1$
BIC $1618.8$ $1619.6$ $1624.3$ $1629.5$
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table O: Tropical cyclones: candidate severity models
A.3.5 Storms
Model 1* Model 2 Model 3
Intercept 11.573-11.573*** 161.061161.061** $71.386$
(<<0.001) ($0.008$) ($0.364$)
SST 0.3480.348** 0.2600.260+
($0.003$) ($0.073$)
MSLP 0.002-0.002** 0.001-0.001
($0.007$) ($0.291$)
AIC $516.1$ $518.2$ $517.0$
BIC $526.5$ $528.6$ $532.6$
Log.Lik. 256.058-256.058 257.101-257.101 255.499-255.499
F $8.800$ $7.240$ $4.888$
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table P: Storms: candidate frequency models
Model 1* Model 2 Model 3
SST $0.239$ $0.176$
($0.103$) ($0.296$)
MSLP 0.001-0.001 0.001-0.001
($0.149$) ($0.471$)
AIC $2504.3$ $2504.9$ $2505.8$
BIC $2510.7$ $2511.3$ $2514.3$
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table Q: Storms: candidate severity models
A.3.6 East Coast Low
Model 1 Model 2*
Intercept 2.593-2.593 2.260-2.260*
($0.518$) ($0.041$)
SST 0.098-0.098
($0.625$)
SST gradients 2.1892.189+
($0.053$)
AIC $81.7$ $77.8$
BIC $90.7$ $86.8$
Log.Lik. 38.828-38.828 36.888-36.888
F $0.239$ $3.732$
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table R: East Coast Low: candidate frequency models
Given the limited data and high uncertainty in projected East Coast Low intensities under climate change (Pepler et al., 2016b), we assume a stationary distribution for the normalized severity of East Coast Lows.
A.3.7 Hailstorms
Model 1
Intercept 11.589-11.589*
($0.014$)
Atmospheric temperature 0.079-0.079
($0.452$)
Near-surface temperature 0.2110.211***
(<<0.001)
AIC $281.3$
BIC $296.9$
Log.Lik. 137.654-137.654
F $6.437$
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table S: Hailstorm: frequency model
As all coefficients are statistically insignificant and their signs do not align with expectations (e.g., a positive relationship between near-surface temperature and hailstorm intensity is anticipated), we adopt a stationary Log-Normal distribution to model hailstorm severity.
Model 1
Atmospheric temperature 0.049-0.049
($0.724$)
Near-surface temperature 0.083-0.083
($0.394$)
AIC $1304.8$
BIC $1310.4$
+ p << 0.1, * p << 0.05, ** p << 0.01, *** p << 0.001
Table T: Hailstorm: severity model
B Model simulation results: supplementary details
B.1 Flood Loss: Investigation of simulation peaks
The simulated flood losses presented in Section 3.2.1 of the main paper indicate a pronounced peak under the SSP 8.5 scenario, particularly at the $95^{\text{th}}$ percentile. As shown in Figure A, the peak in simulated flood losses around the year 2047 is primarily driven by the increase in projected extreme precipitation (the climate covariate used to model flood frequency and severity), which reaches a record-high value during this period. This surge in projected extreme precipitation reflects the inter-annual variability within the CMIP6 ensemble simulations (Jain et al., 2023), as well as the projected rise in the frequency of extreme ENSO events under high-emission scenarios reported in the literature (see, e.g., Cai et al., 2015; Heede and Fedorov, 2023).
Figure A: Simulated extreme precipitation (rx5day) at the 95th percentile under the SSP 8.5 scenario (left), and simulated flood losses at the 95th percentile under the same scenario (right). The red dashed line marks the peak in the simulated flood loss.
B.2 Bushfire Loss: Simulation with extended projection horizon
To better illustrate the long-term trend in bushfire losses, we extend the simulation horizon of normalised bushfire losses to 2100, as shown in Figure B. The results reveal a clear upward trend under the high-emission scenarios SSP 8.5 and SSP 7.0 across both the mean and higher quantiles. Specifically, the mean bushfire loss under SSP 8.5 is projected to increase by $91\%$ by 2100 relative to its initial value, and to be $95\%$ higher than the corresponding mean loss under SSP 2.6 at that time. This increasing trend aligns with the projected rise in fire susceptibility in Australia (Australian Climate Service, 2025, pp. 32–37).
Figure B: Simulation results of normalised bushfire losses extended to 2100. Solid lines represent the average simulation paths under different climate scenarios, while dashed lines indicate the $5^{\text{th}}$ and $95^{\text{th}}$ percentiles.
C Equity models: supplementary details
C.1 Modelling assumptions of the equity model
In Section 2.3.3 of the main paper, equity returns are modelled as a function of operating profit growth, which in turn is modelled as a function of consumption growth. In this section, we examine the validity of this modelling assumption.
Dividend growth is a critical driver of total equity returns (see, e.g., Lettau and Ludvigson, 2005). The total dividends paid by corporations can be expressed as:
$D_{t}=(\beta_{t}(1-\text{Tax}_{t})\zeta_{t})C_{t},$ (C.1)
where $\beta_{t}$ is the share of corporate’s operating profits in the total economy at time $t$, $\text{Tax}_{t}$ is the tax rate, and $\zeta_{t}$ is the dividend payout ratio at time $t$. Consequently, dividend growth is given by:
$\Delta d_{t}=\frac{(\beta_{t}(1-\text{Tax}_{t})\zeta_{t})C_{t}}{(\beta_{t-1}(1-\text{Tax}_{t-1})\zeta_{t-1})C_{t-1}}-1.$ (C.2)
If the share of corporate operating profits, tax rates, and dividend payout ratios remain constant over time, dividend growth would be equal to consumption growth.
However, this assumption may be too restrictive. Instead of assuming a constant dividend payout ratio, we allow its change to be relatively constant over time (i.e., $1+\Delta\zeta_{t}=\zeta_{1}+\epsilon$, with $\epsilon\sim N(0,\sigma^{2})$), then the change in dividend growth could be written as a linear function of the operating profit growth:
$\Delta d_{t}=\zeta_{0}+\zeta_{1}\Delta\text{OP}_{t}+\tilde{\epsilon}_{t},$ (C.3)
where $\Delta\text{OP}_{t}$ is the change in operating profits defined as $\text{OP}_{t}=\beta_{t}(1-\text{Tax}_{t})C_{t}$. This formulation is analogous to our empirical model shown in the main paper.
Furthermore, the change in operating profit can be expressed as a function of consumption growth:
$\Delta\text{OP}_{t}=(1+\Delta\gamma_{t})\Delta C_{t}+\Delta\gamma_{t},$ (C.4)
where $1+\Delta\gamma_{t}=(1+\Delta\beta_{t})(1+\Delta(1-\text{Tax}_{t}))$. Assuming that the changes in the proportion of corporate profits to total consumption (i.e., $\Delta\beta_{t}$) and in tax rates (i.e., $\Delta(1-\text{Tax}_{t})$) are relatively constant over time, the change in operating profit can be modelled as a linear function of consumption growth:
$\Delta\text{OP}_{t}=\alpha_{0}+\alpha_{1}\Delta C_{t}+\tilde{\epsilon}_{t},$ (C.5)
which is consistent with our empirical model shown in the main paper. As shown in Figure A, the assumption of a constant change in tax rates is supported by historical data, with Australian corporate tax rates remaining at $30\%$ since 2002. Moreover, the proportion of corporate earnings is relatively stable over time, as demonstrated in Figure B.
Figure A: Australian corporate tax rates (Data source: Trading Economics (2024)
Figure B: Proportion of aggregate Australian corporate operating profits to total consumption in Australia (Data source: Australian Bureau of Statistics (2024a))
C.2 Simulation results for general and brown investment portfolios
The cumulative equity returns for the general and brown portfolios are presented in Figure C, showing trends broadly consistent with those of the total investment portfolios discussed in Section 3.2.2.
To examine the difference in investment performance between the general and brown portfolios, Figure D plots the relative difference in excess equity returns (i.e., the percentage deviation of the brown portfolio’s excess return from that of the general portfolio). The results are generated under the assumption that, in the absence of transition effects, the brown portfolio yields the same returns as the general portfolio, consistent with the assumption in Gatzert and Özdil (2024).
A positive deviation is projected under the SSP 8.5 scenario throughout the projection horizon, reflecting its assumption of increasing reliance on fossil fuels, which supports operating profit growth in brown energy firms. Under SSP 4.5, positive deviations appear early in the projection period, driven by temporary increases in fossil fuel production, but turn negative later as dependence on fossil fuels slowly declines (O Neill et al., 2017). A similar pattern is observed under SSP 7.0, albeit with smaller excess returns; however, the lower fossil fuel consumption in OECD countries is attributed to the declining population in developed economies assumed in this scenario, rather than to a reduction in fossil fuel dependence (Riahi et al., 2017).
[…]
Figure C: Simulated (log) compounded equity returns on the general portfolio (a) and the brown portfolio (b).
Figure D: Relative difference (in %) between excess equity returns of the brown portfolio and the general portfolio, averaged across all simulation paths.
D Supplementary benchmarking analysis: Comparison of projected damage ratios in literature
The projected economic damage ratios from our climate-dependent DFA model (as presented in Section 3.2.2 of the main paper) are compared with estimates from the Dynamic Integrated Model of Climate and the Economy (DICE), originally proposed by Nordhaus (1992b). The DICE model has been widely adopted in climate economics to estimate the social cost of carbon and evaluate climate policies. For this comparison, we use damage ratio estimates from the two latest versions of the DICE model: DICE-2023 (Barrage and Nordhaus, 2024) and DICE-2016 (Nordhaus, 2018a).
For benchmarking, we focus on three scenarios from the DICE model: baseline, cost-benefit optimal, and 2 $^{\circ}\text{C}$ target. The baseline scenario in DICE assumes a 3.6 $^{\circ}\text{C}$ global temperature increase by 2100 (Barrage and Nordhaus, 2024), which aligns broadly with the SSP 7.0 scenario (4.1 $^{\circ}\text{C}$ by 2100). The cost-benefit optimal scenario projects a 2.6 $^{\circ}\text{C}$ rise by 2100 (Barrage and Nordhaus, 2024), corresponding to SSP 4.5 (2.63 $^{\circ}\text{C}$ by 2100). Lastly, the 2 $^{\circ}\text{C}$ target scenario assumes a 2 $^{\circ}\text{C}$ limit on warming by 2100 (Barrage and Nordhaus, 2024), which is comparable to SSP 2.6 (1.76 $^{\circ}\text{C}$ by 2100) (Riahi et al., 2017). However, caution is warranted, as the narratives underlying the DICE scenarios differ from those of the SSP framework, and here we only focus on temperature alignment. For a detailed discussion of these narratives, see Barrage and Nordhaus (2024).
Figure A presents a comparison between our simulated damage ratios (Section 3.2.2) and the DICE estimates. Since DICE projections are provided at five-year intervals, we apply spline interpolation to convert them to annual values for benchmarking. Our mean projected damage ratios align more closely with DICE-2016 estimates in the early projection horizon but fall below them in later years. Across the entire horizon, our projections are also lower than those from DICE-2023. However, the estimated damage ratios from both DICE-2016 and DICE-2023 generally fall within the 5th–95th percentile prediction intervals of our simulations in the corresponding SSP scenarios, suggesting that our projections encompass their estimated damage paths.
However, caution is warranted when interpreting these benchmarking results. Firstly, the DICE model estimates global climate impacts, whereas our analysis focuses on localised impacts (specifically, Australia). Given the regional heterogeneity in climate risks, economic structures, and infrastructure profiles, differences between the DICE projections and our results are anticipated. Secondly, the DICE model captures the cumulative, long-term economic impacts of global warming using a quadratic damage function (Barrage and Nordhaus, 2024), whereas our simulations focus on the instantaneous impacts of natural catastrophe events, which are more relevant for general insurance applications. We do not account for chronic climate change effects. This distinction may explain the higher damage ratios projected by the DICE model, particularly in the later projection horizon, where cumulative climate impacts on economic outputs become more pronounced.
Figure A: Comparison of projected economic damage ratios: Climate-dependent DFA (solid lines), DICE-2023 (circled markers), and DICE-2016 (triangular markers)
E Model extensions
E.1 Multi-regional Collective Risk Models and portfolio optimisation
The Collective Risk Model used for modelling hazard events, as defined in (2.3) in Section 2.2.2 of the main paper, can be extended to a multi-regional form:
$\tilde{X}_{t}=\sum_{r=1}^{R}\sum_{i=1}^{I}\sum_{m=1}^{M_{t}^{(i,r)}}\tilde{X}_{t}^{(i),m,r},$ (E.1)
where $M_{t}^{(i,r)}$ denotes the number of events of hazard type $i$ in year $t$ occurring in region $r$, and $\tilde{X}_{t}^{(i),m,r}$ represents the insurance loss associated with the $m^{\text{th}}$ event of hazard type $i$ in year $t$ in region $r$. Region-specific climate covariates can then be incorporated into the modelling of hazard frequency and severity.
The aggregated simulated catastrophe (CAT) losses across regions, as defined in (E.1), can be incorporated into the climate-dependent DFA framework to simulate other modelled variables. In particular, the gross catastrophe loss incurred by insurer $j$ can be expressed as:
$\tilde{X}_{t,(j)}=\sum_{r=1}^{R}w_{j}^{(r)}\tilde{X}_{t}^{(r)},\quad\text{where}\quad\tilde{X}_{t}^{(r)}=\sum_{i=1}^{I}\sum_{m=1}^{M_{t}^{(i,r)}}\tilde{X}_{t}^{(i),m,r}.$ (E.2)
Here, $w_{j}^{(r)}$ denotes the exposure of insurer $j$ to region $r$. This allocation can be treated as a decision variable, determined by optimising underwriting performance over a planning horizon under each climate scenario, thereby accounting for the heterogeneous impacts of climate change across regions (Chapter 12 of IPCC, 2021b, pp. 1767–1876). An empirical application of the multi-regional hazard model and portfolio optimisation under the impact of climate change is left for future research.
E.2 Integration of CAT modelling outputs
The modelling framework proposed in this paper is intentionally broad yet simplified. Future research may incorporate more sophisticated component models to enhance precision, such as proprietary catastrophe (CAT) models.
To enable integration into the climate-dependent DFA framework, information on projected hazard frequency and severity under different climate scenarios need to be obtained from catastrophe (CAT) model providers. As leading CAT modelling firms begin to incorporate the impacts of climate change into their models (see, e.g., RMS, 2022), such integration is becoming increasingly feasible.
CAT model outputs, such as those produced by Risk Management Solutions (RMS), typically provide event-level projections of annual hazard frequency and severity. These outputs can be aggregated to parametrise a Collective Risk Model within the DFA framework. The resulting hazard loss projections can then be incorporated into the broader climate-dependent DFA framework to assess the impacts of climate change on general insurers.
F Sensitivity testing of dividend payments
F.1 Modelling of dividend payments
The results on surplus and insolvency probability presented in Section 3.3 are based on the assumption that no dividends are paid from the surplus, consistent with the paper’s scope of providing a baseline model. As a supplementary analysis, this section examines how varying dividend payout ratios, which represent different management strategies, affect the surplus and insolvency probability of general insurers.
Following Taylor (2008), we assume that dividends are distributed as a fixed proportion of the surplus above the target solvency level. In addition, we impose a ceiling on the maximum dividend payment that may be distributed. Under these assumptions, the dividend payment for insurer $j$ is given by:
$D_{t}^{(j)}=\min\left[\max\left(0,\varphi\cdot(K_{t}^{(j)}-K_{t}^{(j),*})\right),\max(0,\Upsilon\cdot\Pi_{t}^{(j)})\right],$ (F.1)
where $K_{t}^{(j),*}$ denotes the target capital level, $\varphi$ is the dividend payout ratio, $\Pi_{t}^{(j)}$ denotes the net profit (i.e., $\Pi_{t}^{(j)}=(1+\tilde{r}_{t}^{(I)})\cdot\tilde{\pi}_{t}^{(j)}+\tilde{r}_{t}^{(I)}\cdot K_{t-1}^{(j)}-(\tilde{X}_{t,(j)}^{\text{net}}+\tilde{X}_{t,(j)}^{\text{NC}})$), and $\Upsilon$ denotes the cap on dividend payments, expressed as a proportion of the net profit. The target capital level is defined as a multiple of the gross written premium: $K_{t}^{(j),*}=S^{*}\pi_{t,(j)}$, where $S^{*}$ represents the target solvency ratio. This formulation ensures that no dividends are paid when the capital level falls below the target solvency threshold, thereby triggering a recovery plan to restore solvency. In addition, the dividend payment is capped at a proportion of net profit (i.e., $\Upsilon$), which is typically determined by the external regulatory environment.
For the sensitivity analysis, we consider dividend payout ratios $\varphi\in\{0.1,0.3,0.5,0.7\}$. The target solvency ratio is set to 3, guided by the average ratio of general insurers’ capital base to gross written premium over a 15-year period (see Figure A) derived from the APRA statistics (APRA, 2024a). The cap on dividend payments is set at 50% of net profit (i.e., $\Upsilon=50\%$), consistent with guidance previously issued by APRA during periods of financial stress (APRA, 2020). We also repeated the sensitivity analysis under a higher ceiling on dividend payments, set at 100% of net profit (i.e., $\Upsilon=100\%$). The results, presented in Section F.3, are generally consistent with those obtained under the default ceiling of 50%.
Figure A: Ratios of capital base to gross written premium for the Australian general insurance market, averaged across market participants (APRA statistics)
F.2 Results of sensitivity testing on dividend payments
The impacts of different dividend policies on the mean surplus and on the differences in mean surplus across scenarios are illustrated in Figure B and Figure C, respectively. As shown in Figure B, the mean surplus decreases as the dividend payout ratio increases, which aligns with expectations since higher dividend payouts slow capital accumulation. Figure C compares the mean surplus across SSP scenarios under each dividend policy. The highest mean surplus remains under the SSP 8.5 scenario, followed by SSP 2.6, SSP 4.5, and SSP 7.0, consistent with the findings in the main paper. However, the difference in mean surplus between SSP 2.6 and SSP 8.5 narrows as the dividend payout ratio increases, which is further supported by the relative differences shown in Table A. Since dividend payments are only triggered when the surplus exceeds the target level, their effects are more pronounced in scenarios with higher surpluses (i.e., SSP 8.5).
The impacts of different dividend policies on insolvency probability and on the differences in insolvency probability across scenarios are shown in Figure D and Figure E, respectively. As shown in Figure D, insolvency probability increases with higher dividend payout ratios across all scenarios, as expected, since slower capital accumulation reduces the buffer available to absorb future losses. The largest increase is observed under the SSP 8.5 scenario, particularly in the later projection years. This supports the argument in Section 3.3 that the low insolvency probability under SSP 8.5, despite its higher physical risk, is primarily driven by faster capital accumulation and stronger investment returns earlier in the projection horizon, which provide greater resilience against later losses. As dividend payouts increase, this capital buffer is diminished, leaving less available capital to absorb future losses. Given that the SSP 8.5 scenario involves the highest physical risk, this reduction in capital resilience leads to the largest rise in insolvency probability among the four scenarios. Figure E compares insolvency probabilities across SSP scenarios, showing that the ranking identified in the main paper (i.e., highest under SSP 7.0 and lowest under SSP 8.5) remains consistent. However, the gap in insolvency probability between SSP 2.6 and SSP 8.5 narrows, reflecting the relatively greater increase in insolvency risk under SSP 8.5, as discussed above.
Figure B: Comparison of mean surplus under different dividend payout ratio assumptions within each SSP scenario.
Figure C: Comparison of mean surplus across SSP scenarios for each dividend payout ratio assumption.
Table A: Average relative difference in mean surplus between SSP 2.6 and SSP 8.5 scenarios under different dividend payout ratios
Dividend payout ratio Average relative difference in mean surplus (SSP 2.6 v.s. SSP 8.5)
0.0 -79.7%
0.1 -58.5%
0.3 -47.6%
0.5 -45.9%
0.7 -45.2%
Figure D: Comparison of insolvency probability under different dividend payout ratio assumptions within each SSP scenario.
Figure E: Comparison of insolvency probability across SSP scenarios for each dividend payout ratio assumption.
Table B: Mean difference in insolvency probability between SSP 2.6 and SSP 8.5 scenarios under different dividend payout ratios
Dividend payout ratios Mean difference in insolvency probability (SSP 2.6 v.s. SSP 8.5)
0.0 0.0341
0.1 0.0274
0.3 0.0217
0.5 0.0206
0.7 0.0196
F.3 Results of sensitivity testing on dividend payments under a higher cap on maximum dividend payments
The following figures present the mean surplus and insolvency probability under different dividend payout assumptions with a 100% net profit cap. Overall, the results are consistent with those reported in the previous section based on the 50% net profit cap.
Figure F: Comparison of mean surplus under different dividend payout ratio assumptions (under 100% net profit cap) within each SSP scenario.
Figure G: Comparison of mean surplus across SSP scenarios for each dividend payout ratio assumption (under 100% net profit cap).
Figure H: Comparison of insolvency probability under different dividend payout ratio assumptions (under 100% net profit cap) within each SSP scenario.
Figure I: Comparison of insolvency probability across SSP scenarios for each dividend payout ratio assumption (under 100% net profit cap).
G Results of sensitivity testing on economic damages
This section investigates the impact of alternative economic damage assumptions on simulated investment returns by calibrating the projected economic damages in this study to those reported in the literature. In particular, we adopt the damage function used in the Dynamic Integrated model of Climate and the Economy (DICE), which has been widely applied in the climate–economy literature since its introduction by Nordhaus (1992b). Specifically, we use the functional form implemented in the latest version, DICE-2023 (Barrage and Nordhaus, 2024):
$\Omega^{*}(t)=\gamma(\Delta T)^{2},$ (G.1)
where $\Omega^{*}(t)$ denotes the ratio of total climate damage to GDP (i.e., the damage ratio), and $\Delta T$ represents the increase in Global Mean Surface Temperature (GMST) relative to the pre-industrial level. The damage parameter $\gamma$ is set to 0.003467 in DICE-2023 (Barrage and Nordhaus, 2024), compared with 0.00227 in the earlier DICE-2016R2 version (Nordhaus, 2018b) and 0.001478 in the original DICE-1992 version (Nordhaus, 1992a). However, as noted in Online Appendix D, caution is warranted since the DICE damage function is calibrated at the global scale, whereas our simulated damages are specific to Australia. Therefore, the benchmarking results presented in this section should be interpreted solely as sensitivity tests illustrating the implications of alternative economic damage assumptions, rather than as predictive estimates.
To derive the damage ratios implied by the DICE damage function under the SSP scenarios, we substitute the projected GMST increases for each SSP pathway into the functional form defined in (G.1). The resulting damage ratios from the DICE-1992, DICE-2016R2 and DICE-2023 versions, together with the damage-adjusted real GDP based on the DICE-2023 damage function, are presented in Figure A. The DICE-2023 parameters are adopted as the benchmark for economic damage, producing more conservative estimates of damage ratios than those under DICE-2016R2 and DICE-1992. For simplicity, we refer to “DICE-implied damage ratios” as those derived from the functional form and parameters of the DICE-2023 damage function in the subsequent sections. As shown in Figure A, although SSP 8.5 yields the highest projected damage ratio, its damage-adjusted real GDP remains the highest across scenarios, reflecting the strong potential economic growth driven by competitive markets and innovation assumed in this pathway (O Neill et al., 2017).
Figure A: Left panel (a) shows the implied damage ratios under SSP scenarios based on the DICE damage functions and parameters from the DICE-2023 (solid lines) and DICE-2016R2 (dashed lines) versions; right panel (b) presents the real GDP projections for Australia under SSP scenarios and the damage-adjusted values based on the DICE-2023 damage function.
To calibrate our projected economic damages with those implied by the DICE-2023 damage function specified in (G.1), we adopt a quantile-mapping approach with linear transformation function (Piani et al., 2010; Qian and Chang, 2021) rather than applying a constant multiplier to scale insured losses to economic damages, as in Section 2.3.3 of the main paper. Specifically, we scale the simulated insurance losses by matching the quantiles of the mean projected total economic damages to those of the DICE-implied damages over the projection horizon:
$Q_{\bm{\Psi^{*}}}(q)=\hat{\alpha}_{q}+\hat{\beta}_{q}Q_{\bm{\Xi}}(q),$ (G.2)
where $\bm{\Psi}^{*}=\{\psi_{1}^{*},\ldots,\psi_{t}^{*},\ldots,\psi_{T}^{*}\}$ denotes the set of DICE-implied economic damages over the projection horizon $T$, with $\psi_{t}^{*}=Y_{t}\Omega^{*}(t)$; and $\bm{\Xi}=\{\mathbb{E}(\tilde{X}_{1}),\ldots,\mathbb{E}(\tilde{X}_{t}),\ldots,\mathbb{E}(\tilde{X}_{T})\}$ denotes the expected insurance damages derived from the hazard module over the same horizon. Here, $Q_{\bm{\Xi}}(q)$ and $Q_{\bm{\Psi^{*}}}(q)$ denote the $q^{\text{th}}$ quantiles of $\bm{\Xi}$ and $\bm{\Psi^{*}}$, respectively, and the parameters $\hat{\alpha}_{q}$ and $\hat{\beta}_{q}$ are estimated via regression. The adjusted simulated economic damage is then given by:
$\psi_{t}=\hat{\alpha}_{q}+\hat{\beta}_{q}\tilde{X}_{t}.$ (G.3)
This adjustment ensures that the magnitudes and ranges of the mean simulated economic damages are broadly consistent with those of the DICE-implied economic damages over the projection horizon. In the following sections, we refer to “DICE-adjusted damage ratios” as the simulated economic damage ratios calibrated to the DICE-implied damage ratios using (G.3).
The mean DICE-adjusted damage ratios based on this approach are illustrated in Figure B. The quantiles of the DICE-adjusted damage ratios are calibrated to the corresponding quantiles of the DICE-implied damage ratios over the defined projection horizon. Despite inter-annual fluctuations in the simulated mean DICE-adjusted damage ratios, driven by the variability in hazard losses discussed in Section 3.2.1 of the main paper, the overall trends remain broadly consistent with the DICE-implied damage ratios across scenarios.
Figure B: Left panel (a) plots the sorted mean DICE-adjusted damage ratios from (G.3) against the sorted DICE-implied damage ratios over the projection horizon to 2060; right panel (b) shows the mean DICE-adjusted damage ratios together with the DICE-implied damage ratios over the same period.
The left panel of Figure C compares the simulated damage-adjusted cumulative economic growth under both the original simulated damage ratios and the DICE-adjusted damage ratios with the damage-unadjusted cumulative growth. While the DICE-adjusted damage ratios lead to a greater reduction in real GDP, the ranking of cumulative economic growth remains consistent across all cases – raw, damage-adjusted, and DICE-adjusted – with SSP 8.5 exhibiting the highest growth, followed by SSP 2.6, SSP 4.5, and SSP 7.0. This ranking aligns with the scenario narratives in O Neill et al. (2017). Although the projected damage ratio is highest under SSP 8.5, the corresponding damage-adjusted GDP remains the largest due to its stronger baseline growth assumptions, which are consistent with the narrative of “the attainment of human development goals, robust economic growth, and highly engineered infrastructure resulting in relatively low challenges to adaptation” under this scenario (O Neill et al., 2017).
The right panel of Figure C shows the cumulative investment returns based on the DICE-adjusted damage ratios, with trends and rankings across scenarios reflecting the corresponding patterns in cumulative economic growth. These results are consistent with those reported in Section 3.2.2, suggesting that investment returns are primarily driven by the underlying potential growth assumptions of each SSP scenario.
Figure C: Left panel (a) shows the raw (log) cumulative economic growth based on real GDP pathways (i.e., unadjusted for climate damages) underlying the SSP scenarios (Riahi et al., 2017) (solid lines), the damage-adjusted (log) cumulative economic growth based on the original simulated damages from the hazard module (dashed lines), and the damage-adjusted (log) cumulative economic growth based on the DICE-adjusted damage ratios (dash-dotted lines). Right panel (b) shows the (log) cumulative total investment returns based on the DICE-adjusted damage ratios, with solid lines representing the mean simulation paths and dashed lines indicating the $5^{\text{th}}$ and $95^{\text{th}}$ percentiles.
It is noted that the DICE damage function represents only one of several approaches to estimating the economic impacts of rising temperatures, and that damage estimates vary substantially across studies, as shown in the left panel of Figure D. The right panel of Figure D compares the damage-adjusted GDP under SSP 2.6 and SSP 8.5 in 2060 with the corresponding baseline GDP assumptions. Despite higher estimated damage ratios under SSP 8.5, the resulting damage-adjusted GDP remains above that under SSP 2.6 across all studies by 2060, owing to the stronger baseline (damage-unadjusted) economic growth assumptions under SSP 8.5. Accordingly, while we do not perform a detailed sensitivity analysis of investment returns to alternative damage specifications, it might be reasonable to expect that the relative trends and ranking of investment returns across scenarios (at least at the mean and median levels) – based on these damage estimates up to 2060 – would remain broadly consistent with the results presented in Section 3.2.2 of the main paper.
Figure D: Left panel shows the projected damage ratios at 2060 under SSP 2.6 and SSP 8.5 from FUND3.8 (Anthoff and Tol, 2014), PAGE09 (Hope, 2011), DICE-2016R2 (Nordhaus, 2018b), DICE-2023 (Barrage and Nordhaus, 2024), and BHM-2015 (Burke et al., 2015). Damage ratios can be interpreted as the percentage decline in GDP relative to the baseline GDP assumptions underlying each climate scenarios, and they do not represent the decline relative to the current GDP level. The right panel presents the corresponding damage-unadjusted and damage-adjusted real GDP at 2060 under SSP 2.6 and SSP 8.5, with adjustments based on the same model-specific damage ratios. The SSP 2.6 projection is not available in BHM-2015 (Burke et al., 2015) and therefore left blank in the figure.